### v2

# Self-embedding similitudes of Bedford-McMullen carpets with dependent ratios

Abstract.
We prove that any non-degenerate Bedford-McMullen carpet does not allow oblique self-embedding similitudes; that is, if $f$ is a similitude sending the carpet into itself, then the image of the $x$-axis under $f$ must be parallel to one of the principal axes. We also establish a logarithmic commensurability result on the contraction ratios of such embeddings. This completes a previous study by Algom and Hochman [Ergod. Th. & Dynam. Sys. 39 (2019), 577–603] on Bedford-McMullen carpets generated by multiplicatively independent exponents, together with a new proof of their non-obliqueness statement.
For the self-similar case, however, we construct a generalized Sierpiński carpet that is symmetric with respect to an appropriate oblique line and hence allows a reflectional oblique self-embedding. As a complement, we prove that if a generalized Sierpiński carpet satisfies the strong separation condition and permits an oblique rotational self-embedding similitude, then the tangent of the rotation angle takes values ±$\pm 1$.
Key words and phrases:
Bedford-McMullen carpets, self-embedding, obliqueness, generalized Sierpiński carpets, logarithmic commensurability, deleted-digit sets.
2010 Mathematics Subject Classification:
Primary 28A80; Secondary 28A78
1. Introduction
One way to gain a better comprehension of the geometric structure of fractal sets is to delve into the study of their self-embedding mappings, especially affine mappings or even similitudes. Note that a similitude $f$ on $\mathbb{R}^{d}$ can be written as $f(x)=\lambda Ox+a$, where $\lambda\geq 0$ is the similarity ratio (also contraction ratio if $\lambda<1$), $O$ is an orthogonal transformation on $\mathbb{R}^{d}$ (the orthogonal part) and $a\in\mathbb{R}^{d}$. Here a rough principle is that the imposition of additional restrictions on these sets usually corresponds to an increase in the constraints placed on their self-embedding mappings.
A pioneering work in this direction is a logarithmic commensurability theorem established by Feng and Wang [10]. This theorem asserts that if $K\subset\mathbb{R}$ is a self-similar set generated by a homogeneous iterated function system (IFS) with the open set condition and $f$ is a similitude with $f(K)\subset K$, then the contraction ratio of $f$ must be a rational power of the common ratio of mappings in the IFS. This result was later extended to higher dimensional cases by Elekes, Keleti and Máthé [7], albeit with the open set condition being substituted by the strong separation condition. If in addition the mappings in the corresponding IFS share a common contraction ratio and orthogonal part, the author [19] proved the relative openness of the embedded image. In [1], Algom characerized all affine self-embeddings of self-similar sets with the strong separation condition.
For self-affine sets, however, there are much fewer existing results. One reason for this disparity is that a natural rescaling approach, which works well in the self-similar case, usually fails in the self-affine case due to the inherent distortion. Nevertheless, the classic family of Bedford-McMullen carpets, with their simple lattice structure and delicate intrinsic properties, presents an optimal subject for investigation. A formal definition is as follows.
Definition 1.1.
Let $n\geq m\geq 2$ be integers and let $\Lambda\subset\{0,1,\ldots,n-1\}\times\{0,1,\ldots,m-1\}$ be non-empty. Set $\varphi_{i,j}$ to be the affine map given by
(1.1) $\varphi_{i,j}(x,y)=\Big{(}\frac{x+i}{n},\frac{y+j}{m}\Big{)},\quad(i,j)\in\Lambda.$
The attractor $K(n,m,\Lambda)$ associated with the self-affine IFS $\{\varphi_{i,j}:(i,j)\in\Lambda\}$ is usually called a Bedford-McMullen carpet when $n>m$, and a generalized Sierpiński carpet when $n=m$. More precisely,
$K(n,m,\Lambda)=\Big{\{}\Big{(}\sum_{k=1}^{\infty}\frac{i_{k}}{n^{k}},\sum_{k=1}^{\infty}\frac{j_{k}}{m^{k}}\Big{)}:(i_{k},j_{k})\in\Lambda\Big{\}}.$
To avoid trivial cases, in this paper we always assume that $1<\#\Lambda<mn$, where ##\## denotes the cardinality.
In [7], Elekes, Keleti and Máthé studied the intersections of Bedford-McMullen carpets with their translated copies. They observed a measure drop phenomenon occurring in these intersections, except in trivial cases. A systematic study on self-embedding similitudes of Bedford-McMullen carpets was conducted by Algom and Hochman [3], who showed that if the horizontal and vertical ratios $n,m$ are multiplicatively independent, then the carpet permits only nearly trivial self-embedding similitudes. To state and extend their findings, let us introduce a convenient terminology.
Definition 1.2 (Obliqueness).
  1. (1)
    A line $L\subset\mathbb{R}^{2}$ is called oblique if it is not parallel to one of the two principle coordinate axes.
  2. (2)
    A similitude $f$ on $\mathbb{R}^{2}$ with positive similarity ratio is called oblique if $f$ sends the $x$-axis to an oblique line.
Algom and Hochman’s main result is the following theorem.
Theorem 1.3 ([3]).
Let $K=K(n,m,\Lambda)$ be a Bedford-McMullen carpet with $\frac{\log n}{\log m}\notin\mathbb{Q}$. Suppose $K$ is not supported on any line and is not a Cartesian product of the unit interval $[0,1]$ and some Cantor set. If $f$ is a similitude sending $K$ into itself, then $f$ is not oblique and has similarity ratio $1$.
The non-degenerate and non-product assumptions on $K$ are natural. For example, if $K=[0,1]\times\{0\}$ or if $K=[0,1]\times C_{1/3}$ (where $C_{1/3}$ denotes the middle third Cantor set), then $f$ clearly need not to be an isometry. Meanwhile, the assumption of the independence between $n$ and $m$ is a key factor preventing oblique and contracting self-embeddings of the carpet. In fact, it is well known that a Cantor $m$-set cannot be embedded into another Cantor $n$-set if $\frac{\log m}{\log n}\notin\mathbb{Q}$, see [9]. Heuristically, this phenomenon rules out oblique embeddings of non-degenerate Bedford-McMullen carpets with independent ratios, and forces the contraction ratio of any embedding similitude to be dependent on $n$ and $m$ simultaneously (so isometry).
Our main purpose is to study to what extent Theorem 1.3 remains true for the dependent case (i.e., $\frac{\log n}{\log m}\in\mathbb{Q}$) or even the self-similar case (i.e., $n=m$). The presence of dependence inherently complicates the self-embedding problem. Fortunately, first we are able to extend the non-obliqueness statement to the dependent case, together with a new proof for the independent case without analyzing the tangent sets of $K$ as in [3].
Theorem 1.4.
Let $K=K(n,m,\Lambda)$ be a Bedford-McMullen carpet that is not supported on any line. If $f$ is a similitude sending $K$ into itself, then $f$ is not oblique.
We remark that the non-obliqueness statement is closely related to the slicing problem of Bedford-McMullen carpets. For example, writing $N$ to be the maximal number of rectangles that are selected in a row in the initial pattern (see Section 2.1 for the definition), it is easy to find some $0\leq y\leq 1$ such that $\operatorname{dim_{H}}(K\cap(\mathbb{R}\times\{y\}))=\frac{\log N}{\log n}$, where $\operatorname{dim_{H}}$ is the Hausdorff dimension. But an oblique self-embedding similitude sends this horizontal slice into an oblique one, and if one can show that any oblique slice of $K$ has Hausdorff dimension strictly less than $\frac{\log N}{\log n}$, then the existence of oblique self-embeddings is denied. A natural question is: when do we have such an upper bound for oblique slices? By Marstrand’s slicing theorem [13], for any oblique line $L$ in $\mathbb{R}^{2}$,
(1.2) $\operatorname{dim_{H}}(K\cap(L+a))\leq\max\{\operatorname{dim_{H}}K-1,0\}\quad\text{for a.e. }a\in L^{\perp}.$
So if $K$ has non-uniform horizontal fibres (that is, there are two rows containing different numbers of selected rectangles), it is easy to check (see (2.7) for the formula of $\operatorname{dim_{H}}K$) that
$\operatorname{dim_{H}}K-1<\frac{1}{\log m}\cdot\log\Big{(}m\cdot N^{\frac{\log m}{\log n}}\Big{)}-1=\frac{\log N}{\log n}.$
So almost all oblique slices have Hausdorff dimension strictly less than $\frac{\log N}{\log n}$. When $\frac{\log n}{\log m}\notin\mathbb{Q}$, it is highly probable that the upper bound specified in (1.2) holds for all oblique slices rather than for only typical ones, which leads to a quick proof for certain special cases of Theorem 1.4. This slicing problem is formally stated in the nice survey [12] and seems quite challenging. See [2, 4] for recent attempts. However, in the dependent case this upper bound may not hold, as indicated in [5].
On the other hand, one cannot hope to extend the isometry statement in Theorem 1.3 to the dependent case. For example, when $n=m^{2}$, it is easy to construct a Bedford-McMullen carpet which is actually a self-similar carpet. In such instances, there exist numerous non-isometric self-embedding similitudes. Nevertheless, we are able to prove a logarithmic commensurability conclusion as follows.
Theorem 1.5.
Let $K=K(n,m,\Lambda)$ be a Bedford-McMullen carpet with $\frac{\log n}{\log m}\in\mathbb{Q}$ and let $f$ be a similitude sending $K$ into itself. If $K$ is not contained in any line, then $\frac{\log\lambda}{\log n}\in\mathbb{Q}$, where $\lambda$ denotes the contraction ratio of $f$.
We outline here some rough ideas on how to prove the above two results. For the non-obliqueness statement, we rescale $f^{k}(K)$ in an appropriate manner so that, in the absence of vacant rows, a geometric observation can lead us to the desired conclusion for both independent and dependent cases. On the other hand, if there is a vacant row, the independent case can be directly addressed using results on the dimensions of projections of Bedford-McMullen carpets as formulated in [11] . For the dependent case, we first rescale $f^{k}(K)$ in a new way so that there is another oblique self-embedding $g$ for which $g(K)\subset C\times\mathbb{R}$, where $C$ is a “generalized” Cantor-$n$ set. Utilizing a nontrivial extension of the logarithmic commensurability theorem in [10], we establish the non-existence of such a $g$, thereby deriving Theorem 1.5.
Finally, we consider the self-similar case when $n=m$, which can be regarded as the “most dependent” case. While one might anticipate a non-oblique statement in this context, it is not hard to to construct a generalized Sierpiński carpet which is symmetric with respect to an appropriate oblique line, thus permitting an oblique reflectional self-embedding. See Example 5.1. So the non-oblique statement fails in the self-similar setting. However, considering rotational self-embeddings in lieu of reflectional ones and under the assumption of the strong separation condition, we indeed eliminate almost all possibilities. Here a generalized Sierpiński carpet $K=K(n,\Lambda)$ is said to satisfy the strong separation condition if elements in $\{\varphi_{i,j}(K)\}_{(i,j)\in\Lambda}$ are mutually disjoint.
Theorem 1.6.
Let $K=K(n,\Lambda)$ be a generalized Sierpiński carpet satisfying the strong separation condition. Let $\theta\in\mathbb{R}$ and let $R_{\theta}$ be the counterclockwise rotation at the origin by angle $\theta$. If $f(x)=\lambda R_{\theta}x+a$ is an oblique similitude sending $K$ into itself, then $|\tan\theta|=1$.
Theorem 1.6 is achieved as follows. Note that the convex hull of any non-degenerate Sierpiński carpet is a polygon. On the one hand, we show that if there exists an oblique rotational $f$, then the rotational angle is a rational multiple of $\pi$, and simultaneously can be expressed as the sum of finitely many inner angles (modulus $2\pi$) of the above polygon. On the other hand, we prove that all of these inner angles have rational tangents. So the rotation angle has a rational tangent, and the desired conclusion follows directly from a theorem of Niven (see Lemma 5.7). Here the strong separation condition plays a crucial role in deriving the logarithmic commensurability of the contraction ratio of $f$ (by a result in [7]). In fact, by employing some mass comparison arguments, the separation condition in Theorem 1.6 could be relaxed to the requirement of total disconnectedness; however, this does not present a substantial improvement and we will not explore it further within the scope of this paper.
Organization
In the subsequent section, we provide some useful observations on Bedford-McMullen carpets and deleted-digit sets. In Section 3, we present our new proof of the non-obliqueness statement in the independent case. Section 4 is devoted to the proofs of Theorems 1.4 and 1.5. Finally, we establish Theorem 1.6 in Section 5.
Notation
We write $\mathcal{S}(\mathbb{R}^{d})$ to be the collection of all similitudes on $\mathbb{R}^{d}$ with positive similarity ratio. For any collection $\mathcal{A}$ of sets in $\mathbb{R}^{d}$, $\bigcup\mathcal{A}$ is the union of sets in $\mathcal{A}$. For $A\subset\mathbb{R}^{d}$ and $x\in\mathbb{R}^{d}$, we denote by $A+x$ (or $x+A$) the translation of $A$ by $x$, that is, $A+x:=\{a+x:a\in A\}$. For any oblique line $L\subset\mathbb{R}^{2}$, $\pi_{L}$ denotes the orthogonal projection onto $L$. On the other hand, we write $\pi_{1}$ and $\pi_{2}$ to be the orthogonal projections onto the $x$-axis and $y$-axis, respectively. Finally, the concatenation $fg$ of any two mappings $f,g$ simply means their composition $f\circ g$.
2. Preliminaries
2.1. Bedford-McMullen carpets
A Bedford-McMullen carpet $K=K(n,m,\Lambda)$ can be obtained by a standard iteration process as follows. One first divides the unit square $[0,1]^{2}$ into an $n\times m$ grid, selecting a subset of rectangles formed by the grid (called the initial pattern) and then repeatedly substituting the initial pattern on each of the selected rectangles. The limit set is just $K$. For $k\geq 1$, we call every element in
$\{\varphi_{i_{1},j_{1}}\circ\cdots\circ\varphi_{i_{k},j_{k}}([0,1]^{2}):(i_{1},j_{1}),\ldots,(i_{k},j_{k})\in\Lambda\}$
a level-$k$ rectangle. For any level-$k$ rectangle $R$, we write $\varphi_{R}$ to be the natural affine map sending $[0,1]^{2}$ to $R$.
We first record two geometric observations of Bedford-McMullen carpets, which rely heavily on the self-affinity (more precisely, the fact that $n>m$) and might be of independent interest.
Proposition 2.1.
Let $K=K(n,m,\Lambda)$ be a Bedford-McMullen carpet. If there is at most one selected rectangle in each row, then for any oblique line $L$, $\pi_{L}(K)$ is an infinite set.
The condition that there is at most one selected rectangle in each row is not necessary, but the above version will suffice for our purpose.
Proof.
It suffices to consider when $L$ passes through the origin. In this case, there is an angle $\theta$ such that $\pi_{L}(x)$ is essentially given by $\pi_{1}(x)\cos\theta+\pi_{2}(x)\sin\theta$. Since $L$ is oblique, $\cos\theta$ and $\sin\theta$ are both nonzero.
Let $p\geq 1$ and pick a large $k$ so that $\frac{m^{k+p}}{n^{k}}<|\frac{\sin\theta}{\cos\theta}|$. Let $R$ be any level-$k$ rectangle. Since there is at most one selected rectangle in each row, there are at least (#$(\#\Lambda)^{p}$ many level-$(k+p)$ rectangles in $R$ that lies in different rows in the $n^{-k-p}\times m^{-k-p}$ grid. Since every level-$(k+p)$ rectangle intersects $K$, we can select (#$(\#\Lambda)^{p}/2$ many points in $\varphi_{R}(K)$ with $m^{-k-p}$-separated $y$-coordinates. Then their projections under $\pi_{L}$ are mutually distinct: otherwise, there are two of these points, say $a,b$, such that $\pi_{L}(a)=\pi_{L}(b)$, which implies that
$\frac{|\sin\theta|}{|\cos\theta|}$ $=\frac{|\pi_{1}(a)-\pi_{1}(b)|}{|\pi_{2}(a)-\pi_{2}(b)|}$ (since $\pi_{L}(a)=\pi_{L}(b)$)
$\leq\frac{n^{-k}}{|\pi_{2}(a)-\pi_{2}(b)|}$ (since $a,b\in R$)
$\leq\frac{n^{-k}}{m^{-k-p}}$ (by the choice of that subset)
$=\frac{m^{k+p}}{n^{k}}<\frac{|\sin\theta|}{|\cos\theta|},$ (by our choice of $k$)
a contradiction. In particular, $\pi_{L}(K)\supset\pi_{L}(\varphi_{R}(K))$ contains at least (#$(\#\Lambda)^{p}/2$ many points. Since #$\#\Lambda\geq 2$ and $p$ is arbitrary, $\pi_{L}(K)$ must be an infinite set. ∎
Proposition 2.2.
Let $K=K(n,m,\Lambda)$ be a Bedford-McMullen carpet. Then $K$ does not contain any oblique segment of positive length.
When $n=m$, this proposition sometimes fails. For example, when $\Lambda=\{(i,i):0\leq i\leq n-1\}$, the carpet is nothing but a diagonal of the unit square.
Proof.
Since #$\#\Lambda<mn$, there are $10\leq i_{0}\leq n-1,0\leq j_{0}\leq m-$ 1 such that $(i_{0},j_{0})\notin\Lambda$. Denote by $Q$ the open rectangle $=(\frac{i_{0}}{n},\frac{i_{0}+1}{n})\times(\frac{j_{0}}{m},\frac{j_{0}+1}{m})$. Since $(i_{0},j_{0})\notin\Lambda$, $Q\cap K=\varnothing$. By the self-affinity of $K$, for any $k\geq 1$ and any level-$k$ rectangle $R$, $\varphi_{R}(Q)\cap K=\varnothing$. Note that if $R$ is of the form $[\frac{p}{n^{k}},\frac{p+1}{n^{k}}]\times[\frac{q}{m^{k}},\frac{q+1}{m^{k}}]$, then
(2.1) $\varphi_{R}(Q)=\Big{(}\frac{p}{n^{k}}+\frac{i_{0}}{n^{k+1}},\frac{p}{n^{k}}+\frac{i_{0}+1}{n^{k+1}}\Big{)}\times\Big{(}\frac{q}{m^{k}}+\frac{j_{0}}{m^{k+1}},\frac{q}{m^{k}}+\frac{j_{0}+1}{m^{k+1}}\Big{)}.$
Suppose $K$ contains an oblique line segment $\ell$ of length $|\ell|>0$. Let $u$ be the slope of $\ell$. Since $\ell$ is oblique, we may assume without loss of generality that $u>0$. Fix $k$ so large that $\frac{m^{k}u}{n^{k}}<\frac{1}{2}$ and that there exists a level-$k$ rectangle $R$ such that $\ell$ meets both the left and right edges of $R$. The existence of such a rectangle can be found at the end of the proof. For convenience, we also assume that $\ell\subset R$ (considering the truncation $\ell\cap R$ instead). In particular, $\ell\subset\varphi_{R}(K)$ and hence $\varphi_{R}^{-1}(\ell)\subset K$. Note that $\varphi_{R}^{-1}(\ell)$ has slope $\frac{m^{k}u}{n^{k}}$.
Pick a large integer $p$ so that $\frac{1}{m^{p}}<\frac{m^{k}u}{n^{k}}$ and $\frac{1}{n^{p+2}}<\frac{1}{m^{p+3}}$. Let $(0,a)$ be the intersection of $\varphi_{R}^{-1}(\ell)$ and the left edge of $\varphi_{R}^{-1}(R)=[0,1]^{2}$. So $\varphi_{R}^{-1}(\ell)$ meets the right edge of $[0,1]^{2}$ at $(1,a+\frac{m^{k}u}{n^{k}})$. Since $\frac{m^{k}u}{n^{k}}>\frac{1}{m^{p}}\geq\frac{2}{m^{p+1}}$, there is some integer $j$ such that $[\frac{j}{m^{p+1}},\frac{j+1}{m^{p+1}}]\subset[a,a+\frac{m^{k}u}{n^{k}}]$. For convenience, write $g(x):=\frac{m^{k}u}{n^{k}}x+a$ to be the linear function whose graph contains $\varphi_{R}^{-1}(\ell)$.
For $0\leq t\leq n^{p+2}$, set $x_{t}:=\frac{t}{n^{p+2}}$ and $y_{t}:=g(x_{t})$. By our choices of $p$ and $k$,
(2.2) $y_{t+1}-y_{t}=g(x_{t+1})-g(x_{t})=\frac{m^{k}u}{n^{k}}\cdot\frac{1}{n^{p+2}}<\frac{1}{2}\cdot\frac{1}{m^{p+3}}.$
Therefore, there must exist some $t_{0}$ such that
(2.3) $[y_{t_{0}},y_{t_{0}+1}]\subset\Big{[}\frac{j}{m^{p+1}}+\frac{j_{0}}{m^{p+3}},\frac{j}{m^{p+1}}+\frac{j_{0}+1}{m^{p+3}}\Big{]}$
because the interval on the right hand side has length $m^{-(p+3)}$. Write
$\widetilde{R}:=[x_{t_{0}},x_{t_{0}+1}]\times\Big{[}\frac{j}{m^{p+1}},\frac{j}{m^{p+1}}+\frac{1}{m^{p+2}}\Big{]}.$
By (2.3), $\varphi_{R}^{-1}(\ell)$ passes the interior of $\widetilde{R}$. So $\widetilde{R}$ is a level-$(p+2)$ rectangle. By (2.1),
$\varphi_{\widetilde{R}}(Q)=\Big{(}x_{t_{0}}+\frac{i_{0}}{n^{p+3}},x_{t_{0}}+\frac{i_{0}+1}{n^{p+3}}\Big{)}\times\Big{(}\frac{j}{m^{p+1}}+\frac{j_{0}}{m^{p+3}},\frac{j}{m^{p+1}}+\frac{j_{0}+1}{m^{p+3}}\Big{)}.$
Letting $x_{*}:=x_{t_{0}}+\frac{i_{0}}{n^{p+3}}+\frac{1}{2n^{p+3}}$, we have by (2.3) that
$g(x_{*})\in[g(x_{t_{0}}),g(x_{t_{0}+1})]=[y_{t_{0}},y_{t_{0}+1}]\subset\Big{[}\frac{j}{m^{p+1}}+\frac{j_{0}}{m^{p+3}},\frac{j}{m^{p+1}}+\frac{j_{0}+1}{m^{p+3}}\Big{]}$
and hence the point $(x_{*},g(x_{*}))\in\varphi_{\widetilde{R}}(Q)$. But from the definition of $g$, this point is also in $\varphi_{R}^{-1}(\ell)\subset K$, which contradicts that $\varphi_{\widetilde{R}}(Q)\cap K=\varnothing$.
The existence of $R$ can be deduced in a similar way we choose $\widetilde{R}$. Let $(a_{1},b_{1})$ be the left endpoint of the original $\ell$ (before the truncation). Let $x^{\prime}_{0}:=\min\{\frac{i}{n^{k}}:\frac{i}{n_{k}}\geq a_{1}\}$ and $y^{\prime}_{0}$ be such that $(x^{\prime}_{0},y^{\prime}_{0})\in\ell$. Then define recursively that
$x^{\prime}_{t}=x^{\prime}_{t-1}+\frac{1}{n^{k}}\text{ and }y^{\prime}_{t}:=y^{\prime}_{t-1}+\frac{u}{n^{k}},\quad 1\leq t\leq\lfloor|\pi_{1}(\ell)|n^{k}\rfloor-1,$
[…]
$R=[x^{\prime}_{t_{0}-1},x^{\prime}_{t_{0}}]\times\Big{[}\frac{j^{\prime}}{m^{k}},\frac{j^{\prime}+1}{m^{k}}\Big{]}.$
In the rest of this paper, we adopt the following notations. For $K=K(n,m,\Lambda)$,
  • $J:=\{0\leq j\leq m-1:\exists i\text{ s.t. }(i,j)\in\Lambda\}$, which collects the digits of non-empty rows in the initial pattern.
  • $I:=\{0\leq i\leq n-1:\exists j\text{ s.t. }(i,j)\in\Lambda\}$, which collects the digits of non-empty columns in the initial pattern.
  • For $10\leq j\leq m-$ 1, $I_{j}:=\{0\leq i\leq n-1:(i,j)\in\Lambda\}$, which collects the digits of selected rectangles in the $j$-th row.
  • For $10\leq i\leq n-$ 1, $J_{i}:=\{0\leq j\leq m-1:(i,j)\in\Lambda\}$, which collects the digits of selected rectangles in the $i$-th column.
  • $N:=\max_{0\leq j\leq m-1}\#I_{j}$.
  • $K_{x}:=\{y:(x,y)\in K\}$ (vertical slice) and $K^{y}:=\{x:(x,y)\in K\}$ (horizontal slice).
Note for every $y\in\pi_{2}(K)$ with a unique expansion $y=\sum_{k=1}^{\infty}y_{k}m^{-k}$, where $y_{k}\in J$,
$K^{y}$ =$=\Big{\{}\sum_{k=1}^{\infty}\frac{x_{k}}{n^{k}}:\Big{(}\sum_{k=1}^{\infty}\frac{x_{k}}{n^{k}},\sum_{k=1}^{\infty}\frac{y_{k}}{m^{k}}\Big{)}\in K\Big{\}}$
=$=\Big{\{}\sum_{k=1}^{\infty}\frac{x_{k}}{n^{k}}:(x_{k},y_{k})\in\Lambda\text{ for all }k\Big{\}}$
(2.4) =$=\Big{\{}\sum_{k=1}^{\infty}\frac{x_{k}}{n^{k}}:x_{k}\in I_{y_{k}}\text{ for all }k\Big{\}}.$
For future use, we write
(2.5) $K^{y}_{p}:=\bigcup\Big{\{}[x,x+n^{-p}]:x\in\Big{\{}\sum_{k=1}^{p}\frac{x_{k}}{n^{k}}:x_{k}\in I_{y_{k}}\text{ for }1\leq k\leq p\Big{\}}\Big{\}}$
and call the interior of every interval in the above union a basic open interval of $K^{y}_{p}$. It is easy to see that $\{K^{y}_{p}\}_{p=1}^{\infty}$ is decreasing, $\bigcap_{p=1}^{\infty}K^{y}_{p}=K^{y}$ and
(2.6) $\operatorname{dim_{H}}K^{y}\leq\underline{\dim}_{\textup{B}}\,K^{y}\leq\liminf_{M\to\infty}\frac{\log(\prod_{k=1}^{M}\#I_{y_{k}})}{-\log n^{M}}\leq\frac{\log N}{\log n},$
where $\underline{\dim}_{\textup{B}}$ denotes the lower box dimension. Furthermore, it is standard to show that $\mathcal{H}^{\log N/\log n}(K^{y})\leq 1<\infty$ for all $y$ and we omit the proof. Here and afterwards, $\mathcal{H}^{s}$ denotes the $s$-dimensional Hausdorff measure.
The following lemma records some basic facts about these horizontal slices.
Lemma 2.3.
Let $K=K(n,m,\Lambda)$ be a Bedford-McMullen carpet. Then for any $y\in\pi_{2}(K)$ with a unique expansion $y=\sum_{k=1}^{\infty}y_{k}m^{-k}$, $\{y_{k}\}_{k=1}^{\infty}\subset J$,
  1. (1)
    For every $p\geq 1$ and every pair of basic open intervals $U,V$ of $K^{y}_{p}$, $K^{y}\cap U$ is simply a translated copy of $K^{y}\cap V$.
  2. (2)
    For every $p,t\geq 1$ and every basic open interval $U$ of $K^{y}_{p}$, $U\cap K^{y}_{p}$ contain $\prod_{k=p+1}^{p+t}\#I_{y_{k}}$ many basic open intervals of $K^{y}_{p+t}$.
In addition, if #$\#I_{y_{k}}\geq 2$ for all large $k$, then the following properties hold.
  1. (3)
    $\inf_{p\geq 1}\min\{\operatorname{diam}(K^{y}\cap U)/|U|:U\text{ is a basic open interval of $K^{y}_{p}$}\}>0$, where $\operatorname{diam}(\cdot)$ is the diameter.
  2. (4)
    Calling an open interval $G$ a gap of $K^{y}_{p}$ if $G$ is a connected component of $U\setminus K^{y}$, where $U$ is some basic open interval of $K^{y}_{p}$, we have
    $\sup_{p\geq 1}\max\{|G|n^{p}:G\text{ is a gap of }K^{y}_{p}\}<1.$
Proof.
Let $p\geq 1$ and let $U$ be any basic open interval of $K^{y}_{p}$. Writing $a$ to be the left endpoint of $U$, we immediately see from (2.4) and (2.5) that
$K^{y}\cap U$ $=K^{y}\cap(a,a+n^{-p})$
=$=\Big{\{}a+\sum_{k=p+1}^{\infty}\frac{z_{k}}{n^{k}}:z_{k}\in I_{y_{k}}\text{ for all }k\Big{\}}\setminus\{a,a+n^{-p}\}$
$=a+\Big{(}\Big{\{}\sum_{k=p+1}^{\infty}\frac{z_{k}}{n^{k}}:z_{k}\in I_{y_{k}}\text{ for all }k\Big{\}}\setminus\{0,n^{-p}\}\Big{)}.$
This proves (1). Also,
$K^{y}_{p+t}\cap U=\bigcup\Big{\{}\Big{(}a+\sum_{k=p+1}^{p+t}\frac{x_{k}}{n^{k}},\Big{(}a+\sum_{k=p+1}^{p+t}\frac{x_{k}}{n^{k}}\Big{)}+n^{-p-t}\Big{)}:x_{k}\in I_{y_{k}}\Big{\}},$
which implies (2).
Next, let $p$ be so large that #$\#I_{y_{k}}\geq 2$ whenever $k\geq p$. Let $G\subset U$ be any gap of $K^{y}_{p}$. By (2), $U$ clearly contains at least $4$ basic open intervals of $K^{y}_{p+2}$. Since these intervals are disjoint and each of them contains some point in $K^{y}$, $\operatorname{diam}(K^{y}\cap U)\geq n^{-p-2}$ (which proves (4)) and $|G|\leq|U|-2n^{-p-2}=(1-2n^{-2})n^{-p}$, which gives (5). ∎
It is also worthy of mentioning that the Hausdorff dimension of Bedford-McMullen carpets was obtained by Bedford [6] and McMullen [14] independently as follows:
(2.7) $\operatorname{dim_{H}}K=\frac{1}{\log m}\cdot\log\Big{(}\sum_{j=0}^{m-1}(\#I_{j})^{\log m/\log n}\Big{)}.$
2.2. Deleted-digit sets
Deleted-digit sets naturally appear in the study of horizontal and vertical slices and projections of Bedford-McMullen carpets.
Definition 2.4 (Deleted-digit set).
Let $n\geq 2$ be an integer. For any non-empty $\mathcal{D}\subset\{0,1,\ldots,n-1\}$, we write $E(n,\mathcal{D})$ to be the self-similar set associated with the IFS $\{\frac{x+i}{n}:i\in\mathcal{D}\}$ on $\mathbb{R}$. In other words,
$E(n,\mathcal{D})=\Big{\{}\sum_{k=1}^{\infty}\varepsilon_{k}n^{-k}:\varepsilon_{k}\in\mathcal{D}\Big{\}}.$
A common example is the middle-third Cantor set, where $n=3$ and $\mathcal{D}=\{0,2\}$. Note that for any deleted-digit set $E(n,\mathcal{D})$, the associated IFS $\{\frac{x+i}{n}:i\in\mathcal{D}\}$ always satisfies the open set condition, see [8]. Since $\frac{[0,1]+i}{n}\subset[0,1]$ for all $n\geq 1$ and all $10\leq i\leq n-$ 1, we can use the unit interval $[0,1]$ for the standard iteration process to get $E(n,\mathcal{D})$. More precisely, letting $E_{0}(n,\mathcal{D}):=[0,1]$ and recursively define
(2.8) $E_{k}(n,\mathcal{D}):=\bigcup_{i\in\mathcal{D}}\frac{E_{k-1}(n,\mathcal{D})+i}{n},\quad k\geq 1,$
we get a decreasing sequence $\{E_{k}(n,\mathcal{D})\}_{k=0}^{\infty}$ such that $\bigcap_{k=0}^{\infty}E_{k}(n,\mathcal{D})=E(n,\mathcal{D})$.
By a renowned theorem of Schief [17], $E(n,\mathcal{D})$ has non-empty interior if and only if $\operatorname{dim_{H}}E(n,\mathcal{D})=1$. In particular, if $\operatorname{dim_{H}}E(n,\mathcal{D})<1$ then it must be totally disconnected. It is easy to check by induction that $E_{k}(n,\mathcal{D})$ is a union of intervals of the form $[\frac{i}{n^{k}},\frac{i+1}{n^{k}}]$, and each pair of such intervals are either adjacent (that is, they share a common endpoint) or disjoint.
Many horizontal and vertical slices of a Bedford-McMullen carpet $K=K(n,m,\Lambda)$ are deleted-digit sets or a finite union of them. For example, if $y=\sum_{k}jm^{-k}$ for some $j\in J$, then we have by (2.4) that $K^{y}=E(n,I_{j})$. Similarly, if $x=\sum_{k=1}^{\infty}in^{-k}$ for some $i\in I$, then $K_{x}=E(m,J_{i})$. On the other hand, the projections of $K$ onto the principal axes are also deleted-digit sets:
$\pi_{1}(K)$ =$=\Big{\{}\sum_{k=1}^{\infty}\frac{x_{k}}{n^{k}}:\Big{(}\sum_{k=1}^{\infty}\frac{x_{k}}{n^{k}},\sum_{k=1}^{\infty}\frac{y_{k}}{m^{k}}\Big{)}\in K\Big{\}}$
=$=\Big{\{}\sum_{k=1}^{\infty}\frac{x_{k}}{n^{k}}:\exists\{y_{k}\}\text{ such that }(x_{k},y_{k})\in\Lambda\Big{\}}$
=$=\Big{\{}\sum_{k=1}^{\infty}\frac{x_{k}}{n^{k}}:x_{k}\in I\text{ for all $k$}\Big{\}}=E(n,I)$
and similarly, $\pi_{2}(K)=E(m,J)$.
Lemma 2.5.
Let $E=E(n,\mathcal{D})$ be a deleted-digit set with $1<\#\mathcal{D}<n$. If $g$ is a similitude sending $E$ into itself, then the similarity ratio of $g$ is a rational power of $n$.
Proof.
This is a special case of [10, Theorem 1.1]. ∎
A generalization of this logarithmic commensurability theorem as follows is one of the main ingredients in our proof of the dependent and self-similar cases. Recall that $N:=\max_{0\leq j\leq m-1}\#I_{j}$.
Proposition 2.6.
Let $K=K(n,m,\Lambda)$ be a Bedford-McMullen carpet with $12\leq N\leq n-$ 1 and let $y\in\pi_{2}(K)$. Write $\alpha:=\frac{\log N}{\log n}$. If $0<\mathcal{H}^{\alpha}(K^{y})<\infty$ and there is a contraction $h\in\mathcal{S}(\mathbb{R})$ sending $K^{y}$ into itself, then the contraction ratio of $h$ is a rational power of $n$.
Some idea of the proof stems from an ongoing work of the author and Rao [16].
Proof.
Without loss of generality, assume that $h$ is orientation-preserving. If $y\in\pi_{2}(K)$ has more than one expansion with coefficients in $J$, $y\in\bigcup_{k\geq 1}\frac{\mathbb{Z}}{m^{k}}\cap[0,1]$. In this case, it is easy to see that $K^{y}$ is a finite union of scaled copies of $K^{0}$ and $K^{1}$, say $K^{y}=\bigcup_{s\in\Omega}K^{0,s}\cup\bigcup_{t\in\Gamma}K^{1,t}$, where $\Omega,\Gamma\subset\mathbb{Z}$ and
$K^{0,s}:=\frac{K^{0}+s}{n^{k}}\quad\text{and}\quad K^{1,s}:=\frac{K^{1}+t}{n^{k}}$
for some large integer $k$. Since $h(K^{y})\subset K^{y}$, by Baire’s theorem, some $K^{0,s}$ or $K^{1,s}$, say $K^{0,s}$, must contain an (relative) interior part of $h(K^{y})$ and hence of $h(K^{0,s})$. So we can find a scaled copy $E$ of $K^{0}$ contracting by a factor $n^{-q}$ for some large integer $q$ such that $h(E)\subset K^{0,s}$. Since $K^{0}=E(n,I_{0})$ is a deleted-digit set with positive Hausdorff dimension, by Lemma 2.5, the contraction ratio of $h$ is a rational power of $n$.
Now suppose $y$ has a unique expansion $y=\sum_{k=1}^{\infty}y_{k}m^{-k}$, $\{y_{k}\}\subset J$. We first prove that #$\#I_{y_{k}}=N$ for all large $k$. Suppose on the contrary that #$\#I_{y_{k}}<N$ for infinitely many $k$. Then for any $p\geq 1$, when $\widetilde{M}$ is large enough, we have
(2.9) $\prod_{k=1}^{\widetilde{M}}\#I_{y_{k}}\leq(N-1)^{p}N^{\widetilde{M}-p}=\Big{(}\frac{N-1}{N}\Big{)}^{p}N^{\widetilde{M}}.$
Since $0<\mathcal{H}^{\alpha}(K^{y})<\infty$, there is a constant $\delta_{0}$ such that $\mathcal{H}^{\alpha}_{\delta}(K^{y})>\mathcal{H}^{\alpha}(K^{y})/2>0$ when $0<\delta\leq\delta_{0}$ (for the definition of $\mathcal{H}^{\alpha}_{\delta}$, see [8]). However, we have by (2.9) that
$\mathcal{H}^{\alpha}_{n^{-\widetilde{M}}}(K^{y})\leq\Big{(}\frac{N-1}{N}\Big{)}^{p}N^{\widetilde{M}}\cdot(n^{-\widetilde{M}})^{\alpha}=\Big{(}\frac{N-1}{N}\Big{)}^{p}<\frac{1}{2}\mathcal{H}^{\alpha}(K^{y})$
as long as $p$ and $\widetilde{M}$ are large enough. This is a contradiction.
Denote by $\rho$ the contraction ratio of $h$. Suppose on the contrary that $\frac{\log\rho}{\log n}\notin\mathbb{Q}$. Let $\varepsilon>0$ be a small fixed constant (will be specified later) and pick positive integers $s,t$ such that $-\varepsilon\leq s\cdot\frac{\log\rho}{-\log n}-t<0$. Equivalently, $n^{-t}<\rho^{s}\leq n^{\varepsilon-t}$.
Since #$\#I_{y_{k}}\leq N\leq n-1$ for all $k$, it is not hard to check that
$M:=\sup_{p\geq 1}\max\Big{\{}|B|n^{p}:B\text{ is a connected component of }K^{y}_{p}\Big{\}}\leq 2N<\infty,$
that is, for all $p$, every connected component of $K^{y}_{p}$ contains at most $2N$ basic open intervals of $K^{y}_{p}$. Pick a large $p_{0}\geq 1$ such that there is a connected component $B$ of $K^{y}_{p_{0}}$ consisting of exactly $M$ basic open intervals of $K^{y}_{p_{0}}$ and #$\#I_{y_{k}}=N$ for all $k\geq p_{0}$. Note that $B$ is a closed interval and $h^{s}(K^{y})\subset K^{y}$. We distinguish two cases.
Case 1: $h^{s}(B)$ intersects exactly one connected component of $K^{y}_{p_{0}+t}$, say $B^{\prime}$. In this case, $h^{s}(B\cap K^{y})\subset B^{\prime}\cap K^{y}$. Since $B$ consists of $M$ basic open intervals of $K^{y}_{p_{0}}$ and their intersections with $K^{y}$ are distinct from one another by only a translation (recall Lemma 2.3(1)), $h^{s}(B\cap K^{y})$ contains a $\rho^{s}n^{-p_{0}}$-separated subset of cardinality $M$. Since $\rho^{s}n^{-p_{0}}>n^{-p_{0}-t}$, every basic open interval of $K^{y}_{p_{0}+t}$ in $B^{\prime}$, which has length $n^{-p_{0}-t}$, contains at most one point in that subset. Thus by the maximality of $M$, $B^{\prime}$ contains exactly $M$ basic open intervals of $K^{y}_{p_{0}+t}$. Recall that $0<\mathcal{H}^{\alpha}(K^{y})<\infty$. Writing $c$ to be the $\alpha$-Hausdorff measure of the intersection of $K^{y}$ and any basic open interval in $K^{y}_{p_{0}+t}$, we see from Lemma 2.3(1) that $0<c<\infty$ and $\mathcal{H}^{\alpha}(B^{\prime}\cap K^{y})=Mc$. On the other hand, by Lemma 2.3(2), each basic open interval in $B$ contains exactly $\prod_{k=p_{0}+1}^{p_{0}+t}\#I_{y_{k}}=N^{t}$ ones in $K^{y}_{p_{0}+t}$. So $\mathcal{H}^{\alpha}(B\cap K^{y})=MN^{t}c$ and
$\mathcal{H}^{\alpha}(h^{s}(B\cap K^{y}))=\rho^{\alpha s}\mathcal{H}^{\alpha}(B\cap K^{y})$ $=\rho^{\alpha s}\cdot MN^{t}c$
$>n^{-\alpha t}MN^{t}c$
(2.10) $=N^{-t}\cdot M\cdot N^{t}c=Mc=\mathcal{H}^{\alpha}(B^{\prime}\cap K^{y}),$
which contradicts that $h^{s}(B\cap K^{y})\subset B^{\prime}\cap K^{y}$.
Case 2: $h^{s}(B)$ intersects at least two connected components of $K^{y}_{p_{0}+t}$. In this case, $h^{s}(B)$ contains a hole between those components. More precisely, there exists some integer $i$ such that $(\frac{i}{n^{p_{0}+t}},\frac{i+1}{n^{p_{0}+t}})\cap K^{y}=\varnothing$ and $(\frac{i}{n^{p_{0}+t}},\frac{i+1}{n^{p_{0}+t}})\subset h^{s}(B)$. Write $V:=(\frac{i}{n^{p_{0}+t}},\frac{i+1}{n^{p_{0}+t}})$. Comparing the length, we can find either exactly one or two adjacent basic open intervals of $K^{y}_{p_{0}}$ in $B$ of which the images under $h^{s}$ meet $V$. If $V\subset h^{s}(\widetilde{I})$ for only one basic open interval $\widetilde{I}$ in $B$ (see Figure 1(A)), then, since $h^{s}(\widetilde{I}\cap K^{y})\subset K^{y}$, $h^{s}(\widetilde{I}\cap K^{y})\cap V=\varnothing$. This in turn tells us that $\widetilde{I}$ contains a gap of length $\geq|h^{-s}(V)|$. Therefore,
(2.11) $\max\{|G|n^{p_{0}}:G\text{ is a gap of }K^{y}_{p_{0}}\}\geq|h^{-s}(V)|n^{p_{0}}=\rho^{-s}n^{-p_{0}-t}\cdot n^{-p_{0}}>n^{-\varepsilon}.$
Choosing $\varepsilon$ small enough at the beginning, this contradicts Lemma 2.3(4).
Now, suppose there are two basic open intervals $\widetilde{I}_{1},\widetilde{I}_{2}$ of $K^{y}_{p_{0}}$ in $B$ such that $V\cap h^{s}(\widetilde{I}_{i})\neq\varnothing$, $1\leq i\leq 2$. Without loss of generality, assume that $\widetilde{I}_{1}$ is to the left of $\widetilde{I}_{2}$. Denote by $a_{i}$ the left endpoint of $h^{s}(\widetilde{I}_{i})$. See Figure 1(B) for an illustration.
$h^{s}(\widetilde{I})$$V$
(a) One-interval case
[…]
(b) Two-intervals case
Figure 1. An illustration of the local behavior of Case 2
By Lemma 2.3(1),
$\big{(}a_{1},a_{1}+|h^{s}(\widetilde{I}_{2})\cap V|\big{)}\cap K^{y}$ $=\big{(}a_{2},a_{2}+|h^{s}(\widetilde{I}_{2})\cap V|\big{)}\cap K^{y}$
$=h^{s}(\widetilde{I}_{2})\cap V\cap K^{y}$
==$\varnothing$.= ∅ .
Thus
$h^{s}(\widetilde{I}_{1})\cap K^{y}\subset(h^{s}(\widetilde{I}_{1})\setminus V)\setminus\big{(}a_{1},a_{1}+|h^{s}(\widetilde{I}_{2})\cap V|\big{)}.$
But the right hand side is an interval of length
(2.12) $|h^{s}(\widetilde{I}_{1})|-|h^{s}(\widetilde{I_{1}})\cap V|-|h^{s}(\widetilde{I}_{2})\cap V|=\rho^{s}n^{-p_{0}}-|V|<n^{\varepsilon-p_{0}-t}-n^{-p_{0}-t}.$
Choosing $\varepsilon$ small enough at the beginning, this contradicts Lemma 2.3(3). ∎
For convenience, for any compact set $A\subset\mathbb{R}$ and $a\in A$, we call $a$ a left (resp. right) isolated point of $A$ if there is some $r>0$ such that $(a-r,a)\cap A=\varnothing$ (resp. $(a,a+r)\cap A=\varnothing$).
Corollary 2.7.
Let $K,N,y,h$ be as in Proposition 2.6. Then there are left and right isolated points $a,b$ of $K^{y}$ such that $h(a),h(b)$ (or $h(b),h(a)$ if $h$ is orientation-reversing) are also left and right isolated points of $K^{y}$. Moreover, we can guarantee that $\pi_{2}(R_{t})=\pi_{2}(R^{\prime}_{t})$ for all large $t$, where $R_{t},R^{\prime}_{t}$ are two level-$t$ rectangles with $(a,y)\in\varphi_{R_{t}}(K)$ and $(b,y)\in\varphi_{R^{\prime}_{t}}(K)$, respectively.
Roughly speaking, the “moreover” part just requires that the two points $(a,y),(b,y)$ lie in the same row at all stages during the iteration process. This is needed for some technical reason.
Proof.
Suppose first that $y$ has a unique expansion with coefficients in $J$. For the first statement, it suffices to prove it when $h$ is orientation-preserving. In fact, if $h$ is orientation-reversing, then $h^{2}$ is orientation-preserving. If $a,h^{2}(a)$ are both left (resp. right) isolated points of $K^{y}$, then $h(a)$ must be a right (resp. left) isolated point and we are done.
Similarly as before, let $\rho$ be the contraction ratio of $h$. By Proposition 2.6, $\rho=n^{-p/q}$ for some positive integers $p,q$. Applying the argument in the proof of Proposition 2.6 to $s=q$, $t=p$ and $\varepsilon=0$, we see that Case 2 in that proof is impossible because: (1) the right hand side of (2.11) now equals $1$, i.e., $\widetilde{I}$ contains a gap of the same length as itself, which contradicts that $\widetilde{I}\cap K^{y}\neq\varnothing$; (2) the right hand side of (2.12) now equals 00, i.e., $|h^{q}(\widetilde{I_{1}}\cap K^{y})|=0$, which contradicts that $\mathcal{H}^{\alpha}(h^{q}(\widetilde{I_{1}}))>0$. On the other hand, for Case 1 in that proof, we now have $\mathcal{H}^{\alpha}(h^{q}(B\cap K^{y}))=\mathcal{H}^{\alpha}(B^{\prime}\cap K^{y})$. We claim that
(2.13) $h^{q}(B\cap K^{y})=B^{\prime}\cap K^{y}.$
Otherwise, since $h^{q}(B\cap K^{y})\subset B^{\prime}\cap K^{y}$, there exists some $x\in B^{\prime}\cap K^{y}\setminus h^{q}(B\cap K^{y})$. So we can find a large $t$ and a basic open interval $W$ of $K^{y}_{t}$ such that $x\in W$ but $W\cap h^{q}(B\cap K^{y})=\varnothing$. However, since $0<\mathcal{H}^{\alpha}(K^{y})<\infty$, we have again by Lemma 2.3(1) that $\mathcal{H}^{\alpha}(W\cap K^{y})>0$. Thus
$\mathcal{H}^{\alpha}(h^{q}(B\cap K^{y}))\leq\mathcal{H}^{\alpha}(B^{\prime}\cap K^{y})-\mathcal{H}^{\alpha}(W\cap K^{y})<\mathcal{H}^{\alpha}(B^{\prime}\cap K^{y}),$
which is a contradiction.
By the claim, $h^{q}$ maps all the left and right isolated points of $B\cap K^{y}$ to the corresponding isolated points of $B^{\prime}\cap K^{y}$, respectively. Pick any pair of them, say $a^{\prime}$ and $b^{\prime}$. It is easy to check by contradiction that $(h^{q})^{-1}(a^{\prime}),(h^{q})^{-1}(b^{\prime})$ are left and right isolated points of $K^{y}$, respectively. This establishes the first requirement.
For the second one, since $y$ has a unique expansion, $R_{t}$ and $R^{\prime}_{t}$ are unique for all $t$. Moreover, $y$ belongs to the interior of both of $\pi_{2}(R_{t})$ and $\pi_{2}(R^{\prime}_{t})$, which implies that $\pi_{2}(R_{t})=\pi_{2}(R^{\prime}_{t})$.
Now, suppose $y$ has more than one expansion. Similarly as in the proof of Proposition 2.6, we can write $K^{y}=\bigcup_{s\in\Omega}K^{0,s}\cup\bigcup_{t\in\Gamma}K^{1,t}$, where $\Omega,\Gamma\subset\mathbb{Z}$ and
$K^{0,s}:=\frac{K^{0}+s}{n^{k}}\quad\text{and}\quad K^{1,s}:=\frac{K^{1}+t}{n^{k}}$
for some large integer $k$. By Baire’s theorem again, we may assume without loss of generality that some $K^{0,s}$ contains an interior part of $h(K^{y})$. In particular, there is a scaled copy of $K^{0,s}$, say $E=\frac{K^{0}+z}{n^{p}}\subset K^{0,s}$, such that $h(E)\subset K^{0,s}$. Note that
$h(E)\subset K^{0,s}\Longleftrightarrow h\Big{(}\frac{K^{0}+z}{n^{p}}\Big{)}\subset\frac{K^{0}+s}{n^{k}}.$
We can easily reformulate this as $g(K^{0})\subset K^{0}$ for some similitude $g$ with contraction ratio $\rho n^{k-p}$. Since 00 has a unique expansion, as in the first part of this proof, there are nondegenerate open intervals $U,V$ such that $g(U\cap K^{0})=V\cap K^{0}\neq\varnothing$ (recall (2.13)). In particular, $g$ maps all the one-sided isolated points of $K^{0}$ in $U$ to those of $K^{0}$ in $V$. Since the inclusion $g(K^{0})\subset K^{0}$ is just a reformulation of $h(E)\subset K^{0,s}$, $h$ maps all one-sided isolated points of $K^{0,s}$ in $U^{\prime}$ (for some $U^{\prime}$) to one-sided isolated points of $K^{0,s}$ in $V^{\prime}$ (for some $V^{\prime}$). It remains to prove that these points are also one-sided isolated points of $K^{y}$.
To get the desired statement, note first that we may pick $V^{\prime}$ properly so that
(2.14) $K^{0,s}\cap V^{\prime}=K^{y}\cap V^{\prime}.$
In fact, if $t\neq s$ for all $t\in\Gamma$, then we already have (2.14). If some $t$ equals $s$, then $K^{0,s}$ and $K^{1,t}=K^{1,s}$ may overlap. However, note that $\operatorname{dim_{H}}K^{0,s}\geq\operatorname{dim_{H}}K^{y}\geq\operatorname{dim_{H}}K^{1,s}$ (recall that $h(K^{y})\subset K^{0,s}$), we have either $K^{1,s}=K^{0,s}$ or these two deleted-digit sets have different patterns (in particular, $K^{0,s}\neq K^{1,s}$). So there is a small open interval $V^{\prime\prime}\subset V^{\prime}$ such that $(K^{0,s}\cap V^{\prime\prime})\cap K^{1,s}=\varnothing$ but $K^{0,s}\cap V^{\prime\prime}\neq\varnothing$. Replacing $V^{\prime}$ with $V^{\prime\prime}$ gives us (2.14). In particular, all left (resp. right ) isolated points of $K^{0,s}$ in $V^{\prime}$ are left (resp. right) isolated points of $K^{y}$.
Next, if a one-sided isolated point $a$ of $K^{0,s}$ in $U^{\prime}$ is not a one-sided isolated point of $K^{y}$, there is a sequence $\{x_{n}\}\subset K^{y}$ such that $x_{n}\to a$ from the left or from the right. Then $h(x_{n})\to h(a)$ from one direction. Since $h(x_{n})\in K^{y}$, this contradicts the fact that $h(a)$ is a one-sided isolated point of $K^{y}$ in $V^{\prime}$, which completes the proof. ∎
3. The independent case
One significant benefit of the independence assumption regarding $n$ and $m$ is that it elucidates the dimensions of the orthogonal projections of $K$ onto all lines.
Lemma 3.1 ([11]).
Let $K=K(n,m,\Lambda)$ be a Bedford-McMullen carpet. If $\frac{\log n}{\log m}\notin\mathbb{Q}$, then $\operatorname{dim_{H}}\pi_{L}(K)=\min\{\operatorname{dim_{H}}K,1\}$ for every oblique line $L$.
Theorem 3.2.
Let $K,f$ be as in Theorem 1.4. If $\frac{\log n}{\log m}\notin\mathbb{Q}$, then $f$ is not oblique.
We distinguish three cases. Recall that $N=\max_{j}\#I_{j}$.
Case 1
#$\#J<m$. In other words, there is an empty row in the initial pattern of $K$.
Proof of Theorem 3.2 under Case 1.
If $f$ is oblique, then by Lemma 3.1,
(3.1) $\operatorname{dim_{H}}\pi_{t}(K)\geq\operatorname{dim_{H}}\pi_{t}(f(K))=\min\{\operatorname{dim_{H}}K,1\},\quad t=1,2.$
If there is some $j$ such that #$\#I_{j}\geq 2$, then
$\min\{1,\operatorname{dim_{H}}K\}$ $>\frac{1}{\log m}\cdot\log(\#J)$ (since #$\#J<m$ and (2.7))
$=\operatorname{dim_{H}}E(m,J)$
$=\operatorname{dim_{H}}\pi_{2}(K),$ (since $E(m,J)=\pi_{2}(K)$)
which contradicts (3.1). However, if #$\#I_{j}=1$ for all $j\in J$, then #$\#\Lambda=\#J$. Therefore,
$\operatorname{dim_{H}}\pi_{1}(K)=\operatorname{dim_{H}}E(n,I)\leq\frac{\log\#\Lambda}{\log n}<\frac{\log\#J}{\log m}=\operatorname{dim_{H}}\pi_{2}(K)\leq\min\{\operatorname{dim_{H}}K,1\},$
which again contradicts (3.1). ∎
Case 2
#$\#J=m$ and $N=n$. In other words, there is no empty row but a full row in the initial pattern of $K$.
Proof of Theorem 3.2 under Case 2.
In this case, pick $j$ such that #$\#I_{j}=n$. Writing $y:=\sum_{k=1}^{\infty}jm^{-k}$, we have seen that $K^{y}=E(n,I_{j})$, which is a horizontal line segment. So if $f$ is oblique, then $K\supset f(K)$ contains an oblique segment, which contradicts Proposition 2.2. ∎
Case 3
#$\#J=m$ but $N<n$. In other words, one can find a selected rectangle and an unselected one in each row in the initial pattern of $K$.
Proof of Theorem 3.2 under Case 3.
[…]
Let us pick a sequence $\{R_{k}\}_{k=1}^{\infty}$, where each $R_{k}$ is a level-$k$ rectangle and $R_{k+1}\subset R_{k}$ for all $k\geq 1$. Write $a_{k}$ to be the left bottom vertex of $R_{k}$ and $Q_{k}$ to be any $n^{-k}\times m^{-k}$ grid rectangle containing $f(a_{k})$. For convenience, we also let $\ell^{k},\ell_{k}$ be the bottom and left edges of $R_{k}$, respectively. So $\ell^{k}\cap\ell_{k}=\{a_{k}\}$. See Figure 2 for an illustration. Below we record two simple but key facts. Here we abuse notation slightly by writing $\varphi_{Q_{k}}$ to be the natural affine map sending $[0,1]^{2}$ onto $Q_{k}$.
  1. (1)
[…]
    $|\varphi_{Q_{k}}^{-1}f(\ell^{k})|$ $=\sqrt{n^{2k}|\pi_{1}(f(\ell^{k}))|^{2}+m^{2k}|\pi_{2}(f(\ell^{k}))|^{2}}$
    $=\sqrt{n^{2k}\cdot\lambda^{2}|\ell^{k}|^{2}\cdot\frac{1}{1+u^{2}}+m^{2k}\cdot\lambda^{2}|\ell^{k}|^{2}\cdot\frac{u^{2}}{1+u^{2}}}$
    $=\sqrt{n^{2k}\cdot\frac{\lambda^{2}}{n^{2k}}\cdot\frac{1}{1+u^{2}}+m^{2k}\cdot\frac{\lambda^{2}}{n^{2k}}\cdot\frac{u^{2}}{1+u^{2}}}$
    (3.2) $\to\frac{\lambda}{\sqrt{1+u^{2}}},\quad k\to\infty$
  2. (2)
    Similarly, the line supporting $\varphi_{Q_{k}}^{-1}f(\ell_{k})$ has slope $-\frac{m^{k}}{n^{k}u}$, which also tends to 00 as $k\to\infty$. Moreover, $\varphi_{Q_{k}}^{-1}f(\ell_{k})$ has length
    $|\varphi_{Q_{k}}^{-1}f(\ell_{k})|$ $=\sqrt{n^{2k}|\pi_{1}(f(\ell_{k}))|^{2}+m^{2k}|\pi_{2}(f(\ell_{k}))|^{2}}$
    $=\sqrt{n^{2k}\cdot\frac{\lambda^{2}}{m^{2k}}\cdot\frac{1}{1+u^{-2}}+m^{2k}\cdot\frac{\lambda^{2}}{m^{2k}}\cdot\frac{u^{-2}}{1+u^{-2}}}$
    $\to\infty,\quad k\to\infty.$
Roughly speaking, for all large $k$, $\varphi_{Q_{k}}^{-1}f(R_{k})$ is a very flat parallelogram with one edge short and the other one pretty long.
[…]
Figure 2. The evolution from $R_{k}$ to $f(R_{k})$ to $\varphi_{Q_{k}}^{-1}f(R_{k})$ (color online)
Write $\widetilde{a}_{k}:=\varphi_{Q_{k}}^{-1}f(a_{k})$. Since $f(a_{k})\in Q_{k}$, $\widetilde{a}_{k}\in[0,1]^{2}$. Fix a large $k$ (will be specified later) so that $\frac{m^{k}}{n^{k}}\cdot\max\{u^{-1},u\}<\lambda$ and $|\varphi_{Q_{k}}^{-1}f(\ell^{k})|\leq\frac{2\lambda}{\sqrt{1+u^{2}}}$ (recall (3.2)). Let us truncate the parallelogram $\varphi_{Q_{k}}^{-1}f(R_{k})$ as $\bigcup_{t\in\mathbb{Z}}\widetilde{E}_{t}$, where
$\widetilde{E}_{t}:=\varphi_{Q_{k}}^{-1}f(R_{k})\cap\{x\in\mathbb{R}^{2}:\pi_{1}(x)\in[-t,-t+1]\}.$
Note that for every $t$,
$|\pi_{2}(\widetilde{E}_{t})|$ $\leq|\pi_{2}(\varphi_{Q_{k}}^{-1}f(R_{k}))|$
$\leq|\pi_{2}(\varphi_{Q_{k}}^{-1}f(\ell^{k}))|+|\pi_{2}(\varphi_{Q_{k}}^{-1}f(\ell_{k}))|$
(3.3) $=m^{k}\cdot\frac{\lambda}{n^{k}}\cdot\frac{u}{\sqrt{1+u^{2}}}+m^{k}\cdot\frac{\lambda}{m^{k}}\cdot\frac{u^{-1}}{\sqrt{1+u^{-2}}}<2\lambda.$
Let us prove Case 3 under the following Condition A first and then explain how to move on without its validity.
Condition A
There are integers $t_{0}\geq-1$ and $0\leq j_{0}\leq m$ and such that $|\pi_{1}(\widetilde{E}_{t_{0}})|=1$ while $\pi_{2}(\widetilde{E}_{t_{0}})\subset[\tfrac{j_{0}}{m},\tfrac{j_{0}+1}{m}]$.
Assume Condition A. Let $j^{\prime}_{0}:=j_{0}\pmod{m}$. Recall that $\varphi_{Q_{k}}^{-1}f(\ell_{k})$ is supported on a line of slope $-\frac{m^{k}}{n^{k}u}$, say $g(x)=-\frac{m^{k}}{n^{k}u}x+c$. Since #$\#I_{j^{\prime}_{0}}<n$, there is some $10\leq i_{0}\leq n-$ 1 such that $K$ does not intersect the open rectangle $(\frac{i_{0}}{n},\frac{i_{0}+1}{n})\times(\frac{j^{\prime}_{0}}{m},\frac{j^{\prime}_{0}+1}{m})=:U$. Write $e$ to be the point $(-t_{0},\lfloor\frac{j_{0}}{m}\rfloor)$ and $x_{0}:=-t_{0}+\frac{i_{0}+1/2}{n}$. Then the point $\xi_{0}:=(x_{0},g(x_{0}))$ lies on the segment $\varphi_{Q_{k}}^{-1}f(\ell_{k})$, and
$U+e$ $=\Big{(}\frac{i_{0}}{n}-t_{0},\frac{i_{0}+1}{n}-t_{0}\Big{)}\times\Big{(}\frac{j^{\prime}_{0}}{m}+\lfloor\frac{j_{0}}{m}\rfloor,\frac{j^{\prime}_{0}+1}{m}+\lfloor\frac{j_{0}}{m}\rfloor\Big{)}$
$=\Big{(}\frac{i_{0}}{n}-t_{0},\frac{i_{0}+1}{n}-t_{0}\Big{)}\times\Big{(}\frac{j_{0}}{m},\frac{j_{0}+1}{m}\Big{)}.$
Since $\pi_{2}(\widetilde{E}_{t_{0}})\subset[\tfrac{j_{0}}{m},\tfrac{j_{0}+1}{m}]$, it is not hard to check that
(3.4) $\xi_{0}+\varphi_{Q_{k}}^{-1}f(\ell^{k})\subset U+e.$
See Figure 3 for an illustration.
[…]
Figure 3. $\varphi_{Q_{k}}^{-1}f(\ell^{k})+\xi_{0}$ is contained in the hole $U+e$ (color online)
Since $\xi_{0}+\varphi_{Q_{k}}^{-1}f(\ell^{k})$ is a translated copy of $\varphi_{Q_{k}}^{-1}f(\ell^{k})$ and is contained in $\varphi_{Q_{k}}^{-1}f(R_{k})$ (the original flat parallelogram), $f^{-1}\varphi_{Q_{k}}(\xi_{0})+\ell^{k}$ is a translated copy of $\ell^{k}$ and is contained in $R_{k}$. Since $\ell^{k}$ is the bottom edge of the level-$k$ rectangle $R_{k}$, $\ell^{k}\cap\varphi_{R_{k}}(K)=\varphi_{R_{k}}(K^{0})$. So there exists some $y_{0}\in[0,1]$ such that
$(f^{-1}\varphi_{Q_{k}}(\xi_{0})+\ell^{k})\cap\varphi_{R_{k}}(K)=\varphi_{R_{k}}(K^{y_{0}})\neq\varnothing,$
where the non-empty conclusion is because $K^{y}\neq\varnothing$ for all $y$ (as pointed out at the beginning of Case 3). This in turn implies that
(3.5) $(\xi_{0}+\varphi_{Q_{k}}^{-1}f(\ell^{k}))\cap\varphi_{Q_{k}}^{-1}f\varphi_{R_{k}}(K)=\varphi_{Q_{k}}^{-1}f\varphi_{R_{k}}(K^{y_{0}})\neq\varnothing.$
Thus
$\varnothing$ $\neq(\xi_{0}+\varphi_{Q_{k}}^{-1}f(\ell^{k}))\cap\varphi_{Q_{k}}^{-1}f\varphi_{R_{k}}(K)$ (just (3.5))
$\subset(U+e)\cap\varphi_{Q_{k}}^{-1}f\varphi_{R_{k}}(K)$ (by (3.4))
$\subset(U+e)\cap\varphi_{Q_{k}}^{-1}f(K)$
$\subset(U+e)\cap\Big{(}\bigcup_{\begin{subarray}{c}\text{$R$ is alevel-$k$ rectangle}\\ \text{$R\cap f(K)\neq\varnothing$}\end{subarray}}\varphi_{Q_{k}}^{-1}\varphi_{R}(K)\Big{)}$ (since $f(K)\subset K$)
$\subset(U+e)\cap\bigcup_{z\in\mathbb{Z}^{2}}(K+z)$
$=(U+e)\cap(K+e)=\varnothing,$ (since $e$ is a lattice point)
which leads to a contradiction.
Now, suppose Condition A is not fulfilled. Set $j_{*}:=\min\{j\in\mathbb{Z}:\pi_{2}(\widetilde{a}_{k})<\frac{j}{m}\}\geq 1$. So the segment $[-2,-1]\times\{\frac{j_{*}}{m}\}$ should intersect $\widetilde{E}_{2}$: otherwise, Condition A holds for $\widetilde{E}_{2}$ and $j_{*}$ or $j_{*}-1$. We claim that
(3.6) $\min\pi_{2}(\widetilde{E}_{2}\cap(\{-2\}\times\mathbb{R}))<\tfrac{j_{*}}{m}<\max\pi_{2}(\widetilde{E}_{2}\cap(\{-1\}\times\mathbb{R})).$
Otherwise, either $\max\pi_{2}(\widetilde{E}_{2}\cap(\{-1\}\times\mathbb{R}))\leq\frac{j_{*}}{m}$ and thus Condition A holds for $\widetilde{E}_{1}$ and $j_{*}-1$; or $\min\pi_{2}(\widetilde{E}_{2}\cap(\{-2\}\times\mathbb{R}))\geq\frac{j_{*}}{m}$ and thus Condition A holds for $\widetilde{E}_{3}$ and $j_{*}$.
If #$\#I_{j^{\prime}}\geq 2$ for some $j^{\prime}$, then we can pick $R_{k}$ properly at the beginning such that
$R_{k}=\varphi_{i_{1},j^{\prime}}\varphi_{i_{2},j^{\prime}}\cdots\varphi_{i_{k},j^{\prime}}([0,1]^{2}),$
where $i_{1},\ldots,i_{k}\in I_{j^{\prime}}$. As a result, selecting $i^{\prime}_{k}\in I_{j^{\prime}}\setminus\{i_{k}\}$, $R^{\prime}_{k}:=\varphi_{i_{1},j^{\prime}}\cdots\varphi_{i_{k-1},j^{\prime}}\varphi_{i^{\prime}_{k},j^{\prime}}([0,1]^{2})$ is another level-$k$ rectangle such that $\pi_{2}(R^{\prime}_{k})=\pi_{2}(R_{k})$ (i.e, they are on the same row of the $n^{-k}\times m^{-k}$ grid) while $\pi_{1}(R^{\prime}_{k})$ is a $\frac{|i^{\prime}_{k}-i_{k}|}{n^{k}}$-translation of $\pi_{1}(R_{k})$. Note that $\varphi_{Q_{k}}^{-1}f(R^{\prime}_{k})$ is simply a translated copy of $\varphi_{Q_{k}}^{-1}f(R_{k})$ along the line supporting $\varphi_{Q_{k}}^{-1}f(\ell^{k})$, and they have disjoint interiors. Picking $k$ large at the beginning (as long as $|\pi_{1}(\widetilde{E}_{2})|=|\pi_{1}(\widetilde{E}^{\prime}_{2})|=1$), the set
$\widetilde{E}^{\prime}_{2}:=\varphi_{Q_{k}}^{-1}f(R^{\prime}_{k})\cap\{x\in\mathbb{R}^{2}:\pi_{1}(x)\in[-2,-1]\},$
is simply a translated copy of $\widetilde{E}_{2}$ (see Figure 4 for an illustration). By (3.6), the segment $[-2,-1]\times\{\frac{j_{*}}{m}\}$ lies entirely in $\widetilde{E}_{2}$ and hence does not meet the interior of $\widetilde{E}^{\prime}_{2}$. Moreover, by the definition of $R^{\prime}_{k}$, it is not hard to check that either (when $\widetilde{E}^{\prime}_{2}$ lies above that segment)
$\frac{j_{*}}{m}\leq\min\pi_{2}(\widetilde{E}^{\prime}_{2})<\max\pi_{2}(\widetilde{E}^{\prime}_{2})\leq\min\pi_{2}(\widetilde{E}_{2})+2n\lambda<\frac{j_{*}+1}{m};$
or (when $\widetilde{E}^{\prime}_{2}$ lies below that segment)
$\frac{j_{*}}{m}\geq\max\pi_{2}(\widetilde{E}^{\prime}_{2})>\min\pi_{2}(\widetilde{E}^{\prime}_{2})\geq\frac{j_{*}}{m}-2n\lambda>\frac{j_{*}-1}{m}.$
Without loss of generality, let us assume the former. So Condition A holds for $\widetilde{E}^{\prime}_{2}$ and $j_{*}$. Applying the previous argument (i.e., the one after assuming Condition A), one can show that the parallelogram $\varphi_{Q_{k}}^{-1}f(R^{\prime}_{k})$ passes through a hole in $\bigcup_{z\in\mathbb{Z}^{2}}K+z$, which leads to a contradiction as before.
[…]
Figure 4. An illustration of the possible positions of $\widetilde{E}^{\prime}_{2}$ (color online)
Finally, if #$\#I_{j}=1$ for all $j$, then since $n>m\geq 2$, we can find $10\leq i_{*}\leq n-$ 1 such that $(i_{*},j_{*}\pmod{m}),(i_{*},j_{*}-1)\notin\Lambda$. Write $j^{\prime}_{*}:=j_{*}\pmod{m}$. Then, since $\pi_{2}(\widetilde{E}_{2})\subset[\tfrac{j_{*}-1}{m},\tfrac{j_{*}+1}{m}]$, applying the previous argument (i.e., the one after assuming Condition A) to $t_{0}=2$, $j_{0}=j_{*}$, $i_{0}=i_{*}$, $j^{\prime}_{0}=j^{\prime}_{*}$ and $U=(\frac{i_{0}}{n},\frac{i_{0}+1}{n})\times(\frac{j^{\prime}_{*}-1}{m},\frac{j^{\prime}_{*}+1}{m})$, one can show that the parallelogram $\varphi_{Q_{k}}^{-1}f(R_{k})$ passes through a hole (which is a translatied copy of $U$) in $\bigcup_{z\in\mathbb{Z}^{2}}K+z$, which leads to a contradiction as before. ∎
Remark 3.3.
Note that in the proof of Cases 2 and 3, we do not need the independence assumption $\frac{\log n}{\log m}\notin\mathbb{Q}$ but only the fact that $n>m$. This will save us much effort in the dependent self-affine case.
4. The dependent self-affine case
In this section, we fix a Bedford-McMullen carpet $K=K(n,m,\Lambda)$ that is not supported in any line and assume that $\frac{\log n}{\log m}\in\mathbb{Q}$. Without loss of generality, write $n=m^{p/q}$, where $p,q\in\mathbb{Z}^{+}$ are coprime. Since $n>m$, $p>q$. Recall that $N=\max_{j}\#I_{j}$.
Unlike the independent case, now we only know the dimension of the projections of $K$ in almost all directions (by Marstrand’s projection theorem). This limitation renders the argument in Case 1 of Theorem 3.2 invalid. Instead, we adopt a rescaling approach as follows.
Lemma 4.1.
Let $f\in\mathcal{S}(\mathbb{R}^{2})$. If $f(K)\subset K$, then we can find $g\in\mathcal{S}(R^{2})$ and a compact set $C\subset\mathbb{R}$ with $\operatorname{dim_{H}}C\leq\frac{\log N}{\log n}$ such that $g(K)\subset C\times\mathbb{R}$. Moreover, if $f$ is oblique, then $g$ can be picked oblique as well.
Proof.
Let $\lambda$ be the contraction ratio of $f$. For all positive integers $t$ with $\lambda^{t}\leq m^{-1}$, write
(4.1) $k_{t}:=\max\{k\geq 0:m^{-pk-1}\geq\lambda^{t}\}.$
Since the contraction ratio of $f^{t}$ is $\lambda^{t}$, $K\subset[0,1]^{2}$ and $f^{t}(K)\subset K$, $f^{t}(K)$ is contained in at most $9$ squares of side length $m^{-pk_{t}}=n^{-qk_{t}}$. More precisely, there are $x\in\frac{\mathbb{N}}{n^{qk_{t}}},y\in\frac{\mathbb{N}}{m^{pk_{t}}}$ such that
$f^{t}(K)\subset(x,x+3n^{-qk_{t}})\times(y,y+3m^{-pk_{t}})=:Q_{t}.$
Denote by $h_{t}$ the homothety sending $(0,1)^{2}$ to the square $(x,x+n^{-qk_{t}})\times(y,y+m^{-pk_{t}})=:Q^{\prime}_{t}$, the “left bottom square” of $Q_{t}$. Note that for all $t$,
$h_{t}^{-1}f^{t}(K)\subset h_{t}^{-1}(Q_{t})=(0,3)^{2}.$
The contraction ratio of $h_{t}^{-1}f^{t}$ equals $n^{qk_{t}}\lambda^{t}$, which is at most $n^{qk_{t}}\cdot m^{-pk_{t}-1}=m^{-1}$ and at least $n^{qk_{t}}\cdot m^{-p(k_{t}+1)-1}=m^{-p-1}$. So the sequence $\{h_{t}^{-1}f^{t}\}_{t}$ has accumulation points in $\mathcal{S}(\mathbb{R}^{2})$. For simplicity, assume that $h_{t}^{-1}f^{t}$ converges to some $g$.
Moreover, write $y=\sum_{k=1}^{pk_{t}}y_{k}m^{-k}$, where $y_{k}\in\{0,\ldots,m-1\}$. Since $p>q$, the collection of all level-$pk_{t}$ rectangles in $\overline{Q^{\prime}_{t}}$, denoted by $\mathcal{R}_{t}$, is (if non-empty)
$\Big{\{}\Big{[}x+\sum_{k=qk_{t}+1}^{pk_{t}}\frac{x_{k}}{n^{k}},\Big{(}x+\sum_{k=qk_{t}+1}^{pk_{t}}\frac{x_{k}}{n^{k}}\Big{)}+n^{-pk_{t}}\Big{]}\times[y,y+m^{-pk_{t}}]:(x_{k},y_{k})\in\Lambda\Big{\}}.$
Then we have
$h_{t}^{-1}f^{t}(K)\cap(0,1)^{2}$ $=h_{t}^{-1}f^{t}(K)\cap h_{t}^{-1}(Q^{\prime}_{t})$
$=h_{t}^{-1}(f^{t}(K)\cap Q^{\prime}_{t})$
(4.2) $\subset h_{t}^{-1}\Big{(}\bigcup\mathcal{R}_{t}\Big{)}$
$\subset\bigcup_{\begin{subarray}{c}(x_{k},y_{k})\in\Lambda,\\ qk_{t}+1\leq k\leq pk_{t}\end{subarray}}\Big{[}\sum_{k=qk_{t}+1}^{pk_{t}}\frac{x_{k}}{n^{k-qk_{t}}},n^{-(p-q)k_{t}}+\sum_{k=qk_{t}+1}^{k_{t}}\frac{x_{k}}{n^{k-qk_{t}}}\Big{]}\times\mathbb{R}$
(4.3) $=:E_{t}\times\mathbb{R}.$
Suppose $g(K)\cap(0,1)^{2}\neq\varnothing$. Then $h_{t}^{-1}f^{t}(K)\cap(0,1)^{2}\neq\varnothing$ for all large $t$. For each of these $t$, note that $\mathcal{R}_{t}\neq\varnothing$ (recall (4.2)) and write $\bm{j}_{t}:=y_{qk_{t}+1}\cdots y_{pk_{t}}$, which is a word of length $(p-q)k_{t}\to\infty$ as $t\to\infty$. Using Cantor’s diagonal argument, it is not hard to find an infinite word $\bm{j}=j_{1}j_{2}\cdots\in J^{\infty}$ such that for all $M\geq 1$, there is a large $t$ such that $|\bm{j}_{t}\wedge\bm{j}|\geq M$, where $\bm{j}_{t}\wedge\bm{j}$ denotes their longest common prefix and ||$\cdot$|| ⋅ | stands for the length. Let $z:=\sum_{k=1}^{\infty}j_{k}m^{-k}$.
We claim that
(4.4) $g(K)\cap(0,1)^{2}\subset K^{z}\times\mathbb{R}.$
If not, then there is some $b\in g(K)\cap(0,1)^{2}$ such that
$\delta:=\operatorname{dist}(\{b\},K^{z}\times\mathbb{R})>0.$
where $\operatorname{dist}(A,B):=\inf\{|a-b|:a\in A,b\in B\}$ is the usual distance. Let $M$ be a large integer with $n^{-M}<\frac{\delta}{4}$. Pick $t$ so large that the following conditions hold:
  1. (1)
    $d_{H}(g(K)\cap(0,1)^{2},h_{t}^{-1}f^{t}(K)\cap(0,1)^{2})<\frac{\delta}{4}$ (where $d_{H}$ is the Hausdorff distance, see [8]);
  2. (2)
    $(p-q)k_{t}\geq M$;
  3. (3)
    there is some $\bm{j}_{t}$ such that $|\bm{j}_{t}\wedge\bm{j}|\geq M$.
The first condition gives us a $b_{1}\in h_{t}^{-1}f^{t}(K)\cap(0,1)^{2}$ with $|b_{1}-b|<\frac{\delta}{4}$. Since $|\bm{j}_{t}\wedge\bm{j}|\geq M$, it is not hard to see from the definition of $E_{t}$ and $\bm{j}_{t}$ that $\operatorname{dist}(E_{t},K^{z})\leq n^{-M}$. Hence
$\operatorname{dist}(\{b_{1}\},K^{z}\times\mathbb{R})$ $\leq\operatorname{dist}(h_{t}^{-1}f^{t}(K)\cap(0,1)^{2},K^{z}\times\mathbb{R})$
$\leq\operatorname{dist}(E_{t},K^{z})$ (by (4.3))
$\leq n^{-M}<\delta/4.$
But it follows that
$\delta=\operatorname{dist}(\{b\},K^{z}\times\mathbb{R})\leq|b-b_{1}|+\operatorname{dist}(\{b_{1}\},K^{z}\times\mathbb{R})<\delta/2,$
which is impossible.
Applying the same argument to the other eight squares in $Q_{t}$, we find at most nine points $z_{0,0}:=z,z_{0,1},\ldots,z_{2,2}\in\pi_{2}(K)$ such that
(4.5) $g(K)\subset\bigcup_{0\leq i,j\leq 2}(K^{z_{i,j}}+i)\times\mathbb{R}.$
Write $C:=\bigcup_{0\leq i,j\leq 2}(K^{z_{i,j}}+i)$. By (2.6), $\operatorname{dim_{H}}C\leq\frac{\log N}{\log n}$.
It remains to prove that if $f$ is oblique then $g$ can be picked oblique as well. Write $O$ to be the orthogonal part of $f$. If $\{O^{t}\}_{t=1}^{\infty}$ is a finite collection, then every element in it appears infinitely many times. Picking an oblique transformation $\widetilde{O}\in\{O^{t}\}_{t}$ and a subsequence $t_{k}$ so that $O^{t_{k}}=\widetilde{O}$ for all $k$, we can consider $f^{t_{k}}$ instead of $f^{t}$ at the beginning of the above arguments and get an oblique $g$ because $g$, as the limit of $h_{t}^{-1}f^{t}$, has $\widetilde{O}$ as its orthogonal part. If $\{O^{t}\}_{t=1}^{\infty}$ is an infinite collection, then since a $2\times 2$ orthogonal matrix is either a rotation or a reflection, $O$ must be an irrational rotation (i.e., the rotation angle is an irrational multiple of $\pi$) and hence every rotation matrix is an accumulation point of $\{O^{t}\}_{t=1}^{\infty}$. Then one just picks a subsequence $\{O^{t_{k}}\}_{k}$ so that $O^{t_{k}}$ converges to an oblique transformation, which, similarly as before, should be the orthogonal part of $g$. This completes the proof. ∎
Remark 4.2.
In the above proof, if we apply Baire’s theorem to (4.5), some $(K^{z_{i,j}}+i)\times\mathbb{R}$ should contain an interior part of $g(K)$. In particular, there is an integer $k$ and a level-$k$ rectangle $R$ such that $g\varphi_{R}(K)\subset(K^{z_{i,j}}+i)\times\mathbb{R}$. As a consequence,
(4.6) $g\varphi_{R}(K^{y}\times\{y\})\subset(K^{z_{i,j}}+i)\times\mathbb{R},\quad\forall y\in\pi_{2}(K).$
We also remark that the integer $k_{t}$ defined in (4.1) equals $\lfloor(-\frac{t\log\lambda}{\log m}-1)/p\rfloor$. So if $\frac{\log\lambda}{\log n}\notin\mathbb{Q}$, then $\frac{\log\lambda}{\log m}\notin\mathbb{Q}$ and hence $\{n^{qk_{t}}\lambda^{t}\}_{t}$ is dense in an open subinterval of $(0,1)$. Since the proof of Lemma 4.1 works for all convergent subsequence of $\{h_{t}^{-1}f^{t}\}_{t=1}^{\infty}$, one can find such an accumulation map $\widetilde{g}$ so that $\widetilde{g}(K)\subset C_{\widetilde{g}}\times\mathbb{R}$, where $C_{\widetilde{g}}$ is a closed set with $\operatorname{dim_{H}}C\leq\frac{\log N}{\log n}$ (just as $C$ in Lemma 4.1), and the contraction ratio of $\widetilde{g}$ is an irrational power of $n$.
Now we are able to prove the non-obliqueness theorem for the dependent case.
Theorem 4.3.
Let $K,m,n$ be as at the beginning of this section. If $f\in\mathcal{S}(\mathbb{R}^{2})$ sends $K$ into itself, then $f$ is not oblique.
Proof.
By Remark 3.3 and Proposition 2.2, it suffices to consider when $N\leq n-1$ and #$\#J\leq m-1$. Suppose on the contrary that $f$ is oblique. Pick $g,C$ as in Lemma 4.1 with $g$ oblique and denote by $\rho$ the contraction ratio of $g$. If $N=1$, then it is not hard to see that every horizontal slice of $K$ is a finite set. From the form of $C$ (recall (4.5)), $C$ is a finite set. But by Proposition 2.1 and the obliqueness of $g$, $\pi_{1}(g(K))$ is an infinite set. This contradicts $\pi_{1}(g(K))\subset\pi_{1}(C\times\mathbb{R})=C$.
Now suppose $N\geq 2$. By Remark 4.2 (see (4.6)), we can find $z\in\pi_{2}(K)$, $c\in\mathbb{R}$, $k\geq 1$ and a level-$k$ rectangle $R$ such that
(4.7) $g\varphi_{R}(K^{y}\times\{y\})\subset(K^{z}+c)\times\mathbb{R},\quad\forall y\in\pi_{2}(K).$
Since $g$ is a similitude, there is a unit directional vector $v:=(v_{1},v_{2})$ such that for every $y\in\pi_{2}(K)$, $g\varphi_{R}$ maps the horizontal line ×$\mathbb{R}\times\{y\}$ to $L_{y}:=\{rv+c_{y}:r\in\mathbb{R}\}$ for some $c_{y}\in\mathbb{R}^{2}$. Together with (4.7), we have
(4.8) $g\varphi_{R}(K^{y}\times\{y\})\subset L_{y}\cap((K^{z}+c)\times\mathbb{R}),\quad\forall y\in\pi_{2}(K).$
Since $g$ is oblique, $v_{1},v_{2}$ are both nonzero. Without loss of generality, assume that $v_{1},v_{2}>0$ (other cases can be similarly discussed).
Note that for each $y\in\pi_{2}(K)$, $g\varphi_{R}(K^{y}\times\{y\})$ can be regarded as a similar copy of $K^{y}$ contracted by $n^{-k}\rho$, while $L_{y}\cap((K^{z}+c)\times\mathbb{R})$ can be regarded as a translated copy of $\frac{|v|}{v_{1}}K^{z}$. In particular, one can obtain from (4.8) a similitude $h_{y}\in\mathcal{S}(\mathbb{R})$ with contraction ratio $n^{-k}\rho\cdot\frac{v_{1}}{|v|}$ (which is independent of $y$) such that
(4.9) $h_{y}(K^{y})\subset K^{z},\quad\forall y\in\pi_{2}(K).$
Pick $j_{*}\in J$ so that #$\#I_{j_{*}}=N$ and let $y_{*}:=\sum_{k=1}^{\infty}j_{*}m^{-k}$. Then $K^{y_{*}}=E(n,I_{j_{*}})$ (recall Section 2.2). By (4.9), $h_{y_{*}}(K^{y_{*}})\subset K^{z}$. So $0<\mathcal{H}^{\alpha}(K^{z})<\infty$, where $\alpha:=\frac{\log N}{\log n}$. It might be suitable for us to point out an observation here (which will be of help later but not in this proof): since (4.9) also implies $h_{z}(K^{z})\subset K^{z}$, it follows from Proposition 2.6 that
(4.10) $\frac{\log(n^{-k}\rho\cdot v_{1}/|v|)}{\log n}\in\mathbb{Q}.$
Next, since $h_{z}(K^{z})\subset K^{z}$, by Corollary 2.7, $h_{z}$ maps a left isolated point $a$ and a right one $b$ both to one-sided isolated points of $K^{z}$. Since $h_{z}(K^{z})\subset K^{z}$ is simply a one-dimensional reformulation of the inclusion $g\varphi_{R}(K^{z}\times\{z\})\subset L_{z}\cap((K^{z}+c)\times\mathbb{R})$, this in turn gives us two points $\widetilde{a},\widetilde{b}\in g\varphi_{R}(K^{z}\times\{z\})$ and some $\delta>0$ such that both of the tubes $(\pi_{1}(\widetilde{a})-\delta,\pi_{1}(\widetilde{a}))\times\mathbb{R}$ and $(\pi_{1}(\widetilde{b}),\pi_{1}(\widetilde{b})+\delta)\times\mathbb{R}$ do not intersect $(K^{z}+c)\times\mathbb{R}$. Moreover, there are two level-$t$ rectangles $R_{t},R^{\prime}_{t}\subset R$ such that $\widetilde{a}\in g\varphi_{R_{t}}(K)$ and $\widetilde{b}\in g\varphi_{R^{\prime}_{t}}(K)$, respectively, and (again by Corollary 2.7) $\pi_{2}(R_{t})=\pi_{2}(R^{\prime}_{t})$ for all $t\geq k_{0}$ for some $k_{0}$. See Figure 5 for an illustration.
[…]
Figure 5. An illustration of $\widetilde{a},\widetilde{b},g(R_{t})$ and $g(R^{\prime}_{t})$, where $(K^{z}+c)\times\mathbb{R}$ is supported in the shaded region. (color online)
Pick $x,x^{\prime}\in K$ with $\pi_{2}(x)=\max\pi_{2}(K)$ and $\pi_{2}(x^{\prime})=\min\pi_{2}(K)$. For convenience, let $e_{1}:=(1,0)$ and $e_{2}:=(0,1)$. Since $\widetilde{a}\in g\varphi_{R_{t}}(K)$, $\widetilde{b}\in g\varphi_{R^{\prime}_{t}}(K)$, $x$ can be expressed as
(4.11) $x=(g\varphi_{R_{t}})^{-1}(\widetilde{a})+c_{t}e_{1}+d_{t}e_{2}=(g\varphi_{R^{\prime}_{t}})^{-1}(\widetilde{b})+\gamma_{t}e_{1}+\eta_{t}e_{2}$
for some $-1\leq c_{t},d_{t},\gamma_{t},\eta_{t}\leq 1$. Since $\pi_{2}(x)$ is the maximum, $d_{t},\eta_{t}\geq 0$. Since $\widetilde{a}$, $\widetilde{b}$ are on the same slice and $\pi_{2}(R_{t})=\pi_{2}(R^{\prime}_{t})$ for $t\geq k_{0}$, we have $d_{t}=\eta_{t}$ for $t\geq k_{0}$. Similarly, there are $-1\leq c^{\prime}_{t},\gamma^{\prime}_{t}\leq 1$ and $-1\leq d^{\prime}_{t}\leq 0$ such that for all $t\geq k_{0}$,
(4.12) $x^{\prime}=(g\varphi_{R_{t}})^{-1}(\widetilde{a})+c^{\prime}_{t}e_{1}+d^{\prime}_{t}e_{2}=(g\varphi_{R^{\prime}_{t}})^{-1}(\widetilde{b})+\gamma^{\prime}_{t}e_{1}+d^{\prime}_{t}e_{2}.$
Fix a large $t\geq k_{0}$ (will be specified later). Note that
$d_{t}-d^{\prime}_{t}=\pi_{2}(x-x^{\prime})=\max\pi_{2}(K)-\min\pi_{2}(K)=:\beta>0.$
So at least one of $d_{t},-d^{\prime}_{t}$ is no less than $\beta/2$. Writing $e_{1,1}:=\pi_{1}(g(e_{1}))$ and $e_{2,1}:=\pi_{1}(g(e_{2}))$, we have by the obliqueness of $g$ that $e_{1,1},e_{2,1}\neq 0$. We distinguish four simple cases.
Case 1: $d_{t}\geq\beta/2$ and $e_{2,1}>0$
In this case, we have by (4.11) that
$g\varphi_{R^{\prime}_{t}}(x)$ $=\widetilde{b}+g\varphi_{R^{\prime}_{t}}(\gamma_{t}e_{1})+g\varphi_{R^{\prime}_{t}}(d_{t}e_{2})$
$=\widetilde{b}+n^{-t}\gamma_{t}g(e_{1})+m^{-t}d_{t}g(e_{2}).$
Since $d_{t}\geq\beta/2$, $n>m$ and $|\gamma_{t}|\leq 1$, $0<n^{-t}\gamma_{t}e_{1,1}+m^{-t}d_{t}e_{2,1}<\delta$ for all large $t$ and hence
$\pi_{1}(g\varphi_{R^{\prime}_{t}}(x))=\pi_{1}(\widetilde{b})+n^{-t}\gamma_{t}e_{1,1}+m^{-t}d_{t}e_{2,1}\in(\pi_{1}(\widetilde{b}),\pi_{1}(\widetilde{b})+\delta).$
In other words, $g\varphi_{R^{\prime}_{t}}(x)$ lies in the tube $(\pi_{1}(\widetilde{b}),\pi_{1}(\widetilde{b})+\delta)\times\mathbb{R}$. But we have seen that this tube does not intersect $(K^{z}+c)\times\mathbb{R}\supset g\varphi_{R}(K)\ni g\varphi_{R^{\prime}_{t}}(x)$, which leads to a contradiction.
Case 2: $d_{t}\geq\beta/2$ and $e_{2,1}<0$
In this case, we have by (4.11) that
$g\varphi_{R_{t}}(x)$ $=\widetilde{a}+g\varphi_{R_{t}}(c_{t}e_{1})+g\varphi_{R_{t}}(d_{t}e_{2})$
$=\widetilde{a}+n^{-t}c_{t}g(e_{1})+m^{-t}d_{t}g(e_{2}).$
Since $d_{t}\geq\beta/2$, $n>m$ and $|c_{t}|\leq 1$, we have $-\delta<n^{-t}c_{t}e_{1,1}+m^{-t}d_{t}e_{2,1}<0$ for all large $t$ and hence
$\pi_{1}(g\varphi_{R_{t}}(x))=\pi_{1}(\widetilde{a})+n^{-t}c_{t}e_{1,1}+m^{-t}d_{t}e_{2,1}\in(\pi_{1}(\widetilde{a})-\delta,\pi_{1}(\widetilde{a})).$
In other words, $g\varphi_{R_{t}}(x)$ lies in the tube $(\pi_{1}(\widetilde{a})-\delta,\pi_{1}(\widetilde{a}))\times\mathbb{R}$. Again, we have seen that this tube does not intersect $(K^{z}+c)\times\mathbb{R}\supset g\varphi_{R}(K)\ni g\varphi_{R_{t}}(x)$, which leads to a contradiction.
Case 3: $d^{\prime}_{t}\geq\beta/2$ and $e_{2,1}>0$
Applying the argument in Case 1 to (4.12), we have for all large $t$ that
$\pi_{1}(g\varphi_{R^{\prime}_{t}}(x^{\prime}))=\pi_{1}(\widetilde{b})+n^{-t}\gamma^{\prime}_{t}e_{1,1}+m^{-t}d^{\prime}_{t}e_{2,1}\in(\pi_{1}(\widetilde{b}),\pi_{1}(\widetilde{b})+\delta),$
which leads to a contradiction.
Case 4: $d^{\prime}_{t}\geq\beta/2$ and $e_{2,1}<0$
Applying the argument in Case 2 to (4.12), we have for all large $t$ that
$\pi_{1}(g\varphi_{R_{t}}(x^{\prime}))=\pi_{1}(\widetilde{a})+n^{-t}c^{\prime}_{t}e_{1,1}+m^{-t}d^{\prime}_{t}e_{2,1}\in(\pi_{1}(\widetilde{a})-\delta,\pi_{1}(\widetilde{a})),$
which leads to a contradiction. ∎
Combining with the above result and Proposition 2.6, we can quickly prove Theorem 1.5.
Proof of Theorem 1.5.
Let $K,n,m$ be as at the beginning of this section. If $f\in\mathcal{S}(\mathbb{R}^{2})$ sends $K$ into itself, then the above theorem tells us that $f$ is not oblique. So it is not hard to find some integer $k\geq 1$ such that $f^{k}(x)=\lambda^{k}x+\eta$ for some $\eta\in\mathbb{R}^{2}$, i.e., the orthogonal part of $f^{k}$ is simply the identity matrix.
Note that $f^{k}$ sends the $x$-axis to a horizontal line. If $\frac{\log\lambda}{\log n}\notin\mathbb{Q}$, we can apply Lemma 4.1 to $f^{k}$ and get (recall the second paragraph in Remark 4.2) some $g\in\mathcal{S}(\mathbb{R}^{2})$ of which the contraction ratio $\rho$ is an irrational power of $n$. However, if $N\geq 2$, applying the argument in the proof of the above theorem (where now $v=(1,0)$) , we can still arrive at (4.10) and hence $\frac{\log\rho}{\log n}\in\mathbb{Q}$. This is a contradiction.
If $N=1$, then every row contains at most one selected rectangle in the initial pattern of $K$. In particular, #$\#\Lambda=\#J\leq m<n$. So there must be a vacant column. On the other hand, since #$\#\Lambda\geq 2$ and $K$ is not contained in any vertical line, #$\#I\geq 2$. Therefore, $\pi_{1}(K)=E(n,I)$ is a deleted-digit set with Hausdorff dimension $10<\frac{\log\#I}{\log n}\leq\frac{\log\#\Lambda}{\log n}<$ 1. Noting that $f^{k}(K)\subset K$ implies that
$\lambda^{k}(E(n,I))+\pi_{1}(\eta)=\lambda^{k}(\pi_{1}(K))+\pi_{1}(\eta)=\pi_{1}(f^{k}(K))\subset\pi_{1}(K)=E(n,I).$
Then it follows from Lemma 2.5 that $\frac{\log\lambda}{\log n}\in\mathbb{Q}$. This completes the proof. ∎
5. The dependent self-similar case
Finally, we consider the case when $m=n$, that is, $K=K(n,\Lambda)$ is a generalized Sierpiński carpet. However, there does exist such a self-similar carpet allowing oblique embeddings.
Example 5.1.
Consider the carpet
$K=\Big{\{}\sum_{k=1}^{\infty}\frac{(x_{k},y_{k})}{4^{k}}:(x_{k},y_{k})\in\{(0,1),(1,3),(2,0),(3,2)\}\Big{\}}.$
See Figure 6. Let $O=\big{(}\begin{smallmatrix}\frac{3}{5}&-\frac{4}{5}\\ -\frac{4}{5}&-\frac{3}{5}\end{smallmatrix}\big{)}$, which is the matrix for the reflection with respect to some line of slope $-1/2$. It is easy to check that
$OK$ $+(\tfrac{9}{5},\tfrac{18}{5})$
=$=\Big{\{}\sum_{k=1}^{\infty}\frac{(\frac{3}{5}x_{k}-\frac{4}{5}y_{k}+\frac{9}{5},-\frac{4}{5}x_{k}-\frac{3}{5}y_{k}+\frac{18}{5})}{4^{k}}:(x_{k},y_{k})\in\{(0,1),(1,3),(2,0),(3,2)\}\Big{\}}$
=$=\Big{\{}\sum_{k=1}^{\infty}\frac{(x_{k},y_{k})}{4^{k}}:(x_{k},y_{k})\in\{(0,1),(1,3),(2,0),(3,2)\}\Big{\}}=K.$
Therefore, $K$ allows an oblique self-embedding $f(x)=Ox+(\frac{9}{5},\frac{18}{5})$.
Figure 6. A carpet allowing oblique self-embeddings
In the rest of this section, we only consider rotational self-embeddings of $K$ and hope to obtain some information on the rotation angle. Recall the notation (1.1). The following result is well known (see e.g. [18]).
Lemma 5.2.
If $K$ is not contained in any line, then the convex hull of $K$ is a polygon. Moreover, every vertex of this polygon is the fixed point of $\varphi_{i,j}$ for some $(i,j)\in\Lambda$.
From now on, let us fix a generalized Sierpiński carpet $K$ satisfying the strong separation condition and not supported in any line. Denote by $P$ the convex hull of $K$. By the above lemma, $P$ is a polygon with finitely many vertices, say $v_{1},\ldots,v_{p}$. For $1\leq t\leq p$, we write $\alpha_{t}$ to be the interior angle of $P$ at the vertex $v_{t}$, and write $\ell_{t}$ to be the edge of $P$ joining $v_{t}$ and $v_{t+1}$. We claim that every $v_{t}$ is not isolated in $K$ even along its adjacent edges.
Corollary 5.3.
For $1\leq t\leq p$, there exists $\{x_{k}\}_{k=1}^{\infty}\subset\ell_{t}\cap K$ such that $x_{k}\to v_{t}$ as $k\to\infty$.
Proof.
Fix any $1\leq t\leq p$. By Lemma 5.2, there is some $(i,j)\in\Lambda$ such that $v_{t}$ is the fixed point of $\varphi_{i,j}$. Then
$v_{t}-\varphi_{i,j}^{k}(v_{t+1})=\varphi_{i,j}^{k}(v_{t})-\varphi_{i,j}^{k}(v_{t+1})=n^{-k}(v_{t}-v_{t+1}),\quad\forall k\geq 1.$
In particular, $\varphi_{i,j}^{k}(v_{t+1})$ must lie on $\ell_{t}$ and tends to $v_{t}$ as $k\to\infty$. ∎
Corollary 5.4.
For each $1\leq t\leq p$, $\tan\alpha_{t}\in\mathbb{Q}$.
Proof.
For $1\leq t\leq p$, we again find by Lemma 5.2 some $(i_{t},j_{t})\in\Lambda$ such that $v_{t}$ is the fixed point of $\varphi_{i_{t},j_{t}}$. Therefore, $v_{t}=(\frac{i_{t}}{n-1},\frac{j_{t}}{n-1})$. This implies that the angle made by every edge $\ell_{t}$ and the $x$-axis has a rational tangent. It follows directly that $\tan\alpha_{t}\in\mathbb{Q}$ for all $t$. ∎
A “local openness” result as follows will be of great help.
Lemma 5.5 ([7, Theorems 4.5 and 4.9]).
Let $f\in\mathcal{S}(\mathbb{R}^{2})$. If $f(K)\subset K$, then there is an open set $U\subset\mathbb{R}^{2}$ such that $U\cap f(K)=U\cap K\neq\varnothing$. Moreover, $\{O^{t}\}_{t=1}^{\infty}$ is a finite collection, where $O$ is the orthogonal part of $f$.
Proposition 5.6.
If $f\in\mathcal{S}(\mathbb{R}^{2})$ sends $K$ into itself, then for every $1\leq t\leq p$, we can find $t^{\prime}$ and two squares $R,R^{\prime}$ (of level $k$ and $k^{\prime}$, respectively) such that the following properties hold:
  1. (1)
    $f\varphi_{R^{\prime}}(v_{t^{\prime}})=\varphi_{R}(v_{t})$,
  2. (2)
    either $f\varphi_{R^{\prime}}(\ell_{t^{\prime}})\subset\varphi_{R}(\ell_{t})$ or $f\varphi_{R^{\prime}}(\ell_{t^{\prime}+1})\subset\varphi_{R}(\ell_{t})$.
Proof.
Fix $1\leq t\leq p$. Without loss of generality, assume that $v_{t}$ lies in the “left bottom area” of $[0,1]^{2}$, that is, $v_{t}\in[0,1/2]^{2}$ (other cases can be similarly discussed). Let us pick an integer $k$ and a level-$k$ square $R=[\frac{i}{n^{k}},\frac{i+1}{n^{k}}]\times[\frac{j}{n^{k}},\frac{j+1}{n^{k}}]$ such that $\varphi_{R}(K)\subset f(K)$ and
$K\cap\big{(}(\tfrac{i-1}{n^{k}},\tfrac{i}{n^{k}})\times(\tfrac{j-1}{n^{k}},\tfrac{j+1}{n^{k}})\big{)}=\varnothing\quad\text{and}\quad K\cap\big{(}(\tfrac{i}{n^{k}},\tfrac{i+1}{n^{k}})\times(\tfrac{j-1}{n^{k}},\tfrac{j}{n^{k}})\big{)}=\varnothing.$
The existence of $R$ will be proved at the end of this proof. As a consequence, there is a small $r>0$ such that
(5.1) $K\cap B(\varphi_{R}(v_{t}),r)=\varphi_{R}(K)\cap B(\varphi_{R}(v_{t}),r)=K\cap\varphi_{R}(P)\cap B(\varphi_{R}(v_{t}),r),$
where $B(x,r)$ denotes the ball centered at $x$ and of radius $r$. Since $\varphi_{R}(v_{t})\in\varphi_{R}(K)\subset f(K)$, there is a level-$k^{\prime}$ square $R^{\prime}$ (for some large $k^{\prime}$) such that $\varphi_{R}(v_{t})\in f\varphi_{R^{\prime}}(K)$. We claim that $\varphi_{R}(v_{t})$ is a vertex of the polygon $f\varphi_{R^{\prime}}(P)$ if $k^{\prime}$ is picked large. In particular, there exists $t^{\prime}$ such that $\varphi_{R}(v_{t})=f\varphi_{R^{\prime}}(v_{t^{\prime}})$, which establishes (1).
To see the claim, note that since $P$ is the convex hull of $K$, $\varphi_{R}(v_{t})\in f\varphi_{R^{\prime}}(K)\subset f\varphi_{R^{\prime}}(P)$. So if $k^{\prime}$ is large enough, then
(5.2) $f\varphi_{R^{\prime}}(P)\subset B(\varphi_{R}(v_{t}),r).$
Therefore,
$f\varphi_{R^{\prime}}(K)$ $\subset f(K)\cap f\varphi_{R^{\prime}}(P)$ (since $K\subset P$)
$\subset K\cap f\varphi_{R^{\prime}}(P)$ (since $f(K)\subset K$)
$\subset K\cap B(\varphi_{R}(v_{t}),r)$ (by (5.2))
[…]
So $f\varphi_{R^{\prime}}(P)\subset\varphi_{R}(P)$. Since $f\varphi_{R^{\prime}}(P)$, $\varphi_{R}(P)$ are both polygons and $\varphi_{R}(v_{t})\in f\varphi_{R^{\prime}}(P)$, $\varphi_{R}(v_{t})$ must be a vertex of $f\varphi_{R^{\prime}}(P)$.
If (2) is false, we see by (1) that $\varphi_{R}(\ell_{t})$ meets $f\varphi_{R^{\prime}}(P)$ only at the vertex $\varphi_{R}(v_{t})$. By Corollary 5.3, there is a sequence $\{x_{p}\}_{p=1}^{\infty}\subset\ell_{t}\cap K$ such that $x_{p}\to v_{t}$ as $p\to\infty$. So $\varphi_{R}(x_{p})\to\varphi_{R}(v_{t})\in f\varphi_{R^{\prime}}(K)$ and $\varphi_{R}(x_{p})\in\varphi_{R}(\ell_{t})$. Since $\varphi_{R}(\ell_{t})$ meets $f\varphi_{R^{\prime}}(P)$ only at the vertex $\varphi_{R}(v_{t})$, $\varphi_{R}(x_{p})\notin f\varphi_{R^{\prime}}(K)$. But recalling that $\varphi_{R}(K)\subset f(K)$, we have $\varphi_{R}(x_{p})\in f(K)$ for all $p$. This implies that
$\operatorname{dist}\big{(}f\varphi_{R^{\prime}}(K),f(K)\setminus f\varphi_{R^{\prime}}(K)\big{)}\leq\inf_{p}\operatorname{dist}(f\varphi_{R^{\prime}}(K),\{\varphi_{R}(x_{p})\})=0,$
which contradicts the strong separation condition.
It remains to prove the existence of $R$. By Lemma 5.5, there is some $k_{0}$ and a level-$k_{0}$ square $Q$ such that $\varphi_{Q}(K)\subset f(K)$. Since $K$ is not contained in any line, we have #$\#I\geq 2$ and #$\#J\geq 2$ (here we abuse slightly these notations from Section 2.1). Iterating the initial pattern if necessary, we may assume that there are $(i_{1},j_{1}),(i_{2},j_{2}),(i_{3},j_{3})\in\Lambda$ such that
$(i_{1},j_{1}-1),(i_{2}-1,j_{2}),(i_{3}-1,j_{3}-1)\in\{0,1,\ldots,n-1\}^{2}\setminus\Lambda.$
Also, write $\underline{i}:=\min I$ and $\underline{j}:=\min J$.
If there is some $11\leq i\leq n-$ 1 such that $(i,\underline{j})\in\Lambda$ but $(i-1,\underline{j})\notin\Lambda$, then it suffices to take $R=\varphi_{Q}\varphi_{i_{1},j_{1}}\varphi_{i,\underline{j}}([0,1]^{2})$. If such a digit does not exist, then $(0,\underline{j})\in\Lambda$. If $I_{\underline{j}}\supsetneq\{0\}$, then $12\leq\#I_{\underline{j}}\leq n-$ 1, where the upper bound holds because $K$ does not contain horizontal segments. So there are at least $4$ level-$2$ squares in the bottom row and it is not hard to see that at least one of them, say $Q^{\prime}$, satisfies that the grid square of side $n^{-2}$ left adjacent to $Q^{\prime}$ is contained in $[0,1]^{2}$ but unselected. Then it suffices to take $R=\varphi_{Q}\varphi_{i_{1},j_{1}}(Q^{\prime})$. So we may assume that $I_{\underline{j}}=\{0\}$. Similarly (looking at $(i_{2},j_{2})$ instead), it suffices to consider when $J_{\underline{i}}=\{0\}$.
Since $(0,\underline{j})\in\Lambda$, $\underline{i}=0$. So $(0,0)\in\Lambda$ but $\{(0,j),(j,0):j\geq 1\}\cap\Lambda=\varnothing$. Then it is easy to see that taking $R=\varphi_{Q}\varphi_{i_{3},j_{3}}\varphi_{0,0}([0,1]^{2})$ would work. ∎
Lemma 5.7 (Niven’s theorem).
If $\theta\in\pi\mathbb{Q}$ and $\tan\theta\in\mathbb{Q}$, then $\tan\theta\in\{0,\pm 1\}$.
Proof.
For a proof, see [15]. ∎
Corollary 5.8.
Let $f(x)=\lambda R_{\theta}x+a$ be as in Theorem 1.6. If $\frac{\theta}{\pi}\in\mathbb{Q}$ and $f$ is an oblique self-embedding similitude of $K$, then $|\tan\theta|=1$.
Proof.
By Lemma 5.7 and the obliqueness of $f$, it suffices to show that $\tan\theta\in\mathbb{Q}$. Fix any $1\leq t\leq p$ and pick $R,R^{\prime},k,k^{\prime},t^{\prime}$ accordingly as in Proposition 5.6. Without loss of generality, assume that $f\varphi_{R^{\prime}}(\ell_{t^{\prime}})\subset\varphi_{R}(\ell_{t})$, which is the first case in Proposition 5.6(2).
$\ell_{t^{\prime}}$$\ell_{t}$$v_{t^{\prime}}$$v_{t}$$\alpha_{t^{\prime}}$$\alpha_{t^{\prime}+1}$$\alpha_{t}$
Figure 7. The rotation effect of $f$ (color online)
Since $\varphi_{R}(\ell_{t})$ is parallel to $\varphi_{R^{\prime}}(\ell_{t})$ and $\varphi_{R^{\prime}},\varphi_{R}$ involve no rotations nor reflections, $f$ sends a line parallel to $\ell_{t^{\prime}}$ to another line parallel to $\ell_{t}$. Without loss of generality, assume that $t^{\prime}<t$ and $v_{1},\ldots,v_{p}$ are in counterclockwise order. Then a simple geometric observation (see Figure 7) tells us that $R_{\theta}$ can be realized as follows: first rotate the line containing $\ell_{t^{\prime}}$ to the one containing $\ell_{t^{\prime}+1}$, and then to the one containing $\ell_{t^{\prime}+2}$ and so on until to $\ell_{t}$. Thus
$\theta=\sum_{p=1}^{t-t^{\prime}}(\pi-\alpha_{t^{\prime}+p})\pmod{2\pi}.$
By Corollary 5.4, $\tan\theta\in\mathbb{Q}$ as desired. ∎
With all these observations in hand, Theorem 1.6 can be easily proved.
Proof of Theorem 1.6.
Again, we may assume that $K$ is not supported in any line. By Lemma 5.5, $\frac{\theta}{\pi}\in\mathbb{Q}$. Then the statement follows directly from the above corollary. ∎
Acknowledgements. I thank Professor Huo-Jun Ruan for reading the manuscript carefully and pointing out several mistakes and a number of typos in an early version of it.